### v1

# AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistency Multispectral CT

Abstract
For nonlinear multispectral computed tomography (CT), accurate and fast image reconstruction is challenging when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched. Motivated by this, we propose an accurate and fast algorithm named AFIRE to address such problem in the case of mildly full scan. We discover that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix. Based on the insights, the AFIRE is proposed respectively from the continuous, discrete and actual-use perspectives by leveraging the simplified Newton method. Under proper conditions, we establish the convergence theory of the proposed algorithm. Furthermore, numerical experiments are also carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in completely geometric-inconsistency dual-energy CT with noiseless and noisy projection data. Particularly, the proposed algorithm significantly outperforms some state-of-the-art methods in terms of accuracy and efficiency. Finally, the flexibility and extensibility of the proposed algorithm are also demonstrated.
keywords:
Multispectral CT, dual-energy CT, geometric inconsistency, large-scale nonlinear inverse problem, accurate and fast image reconstruction algorithm, one-step method, convergence analysis
1 Introduction
Basically, multispectral computed tomography (MSCT) is an imaging modality aiming to reconstruct material-specific (basis) images from multiple sets of attenuation data collected with different X-ray energy spectra, which can subsequently be used to form virtual monochromatic images (VMIs) or the spatial distributions of linear attenuation coefficient (LAC) at energies of interest by linear combination [18, 3, 27]. MSCT has attracted increasing attention and research interest due to the development of hardware and the requirement of applications, as well as its advantages in considering the X-ray energy-dependent information over the conventional CT such as artifact suppression, contrast improvement and object separation [27]. Dual-energy CT (DECT), as the first generation of MSCT, has been implemented in clinical practice for about twenty years [17]. Despite the long history of DECT, its clinical and technical research is still in its infancy [19].
It is well-known that the forward data model used in conventional CT assumes that all of the X-ray photons emitting from an X-ray source are monochromatic, namely, have the same energy [32]. This assumption makes the model linear, which causes the image reconstruction to become a linear inverse problem, but leading to difficulties in artifact suppression and limitations in material differentiation. In real-world CT, however, this assumption is not realistic, since the output X-ray beams are polychromatic, which consist of photons with a wide range of energies that distributed as the energy spectrum. The LAC of the scanned material depends not only on its own elemental composition but also on the energy of the probing X-ray photons. Hence, the forward data model of MSCT may be described as a nonlinear mapping, and the corresponding image reconstruction can be formulated as a nonlinear inverse problem [3, 26, 29, 19, 27, 21].
Although development of algorithms for reconstructing basis images in MSCT constitutes an important topic of active research, accurate and fast image reconstruction remains a challenging task, largely because its forward data model is nonlinear. In this work, we understand different reconstruction algorithms from the perspective of whether the scanning geometric parameters are consistent or matched under different X-ray energy spectra.
One situation is that the scanning geometric parameters are consistent under different energy spectra. That is to say, for each X-ray path, under every energy spectrum, there is always an X-ray beam passing through it. Here this case is called geometric-consistency MSCT, in which the image reconstruction can be accurately realized by a two-step method: step (i) estimates the basis sinograms from measurements through solving a lot of very small-scale independent systems of nonlinear equations, whereas step (ii) reconstructs basis images from the basis sinograms estimated. This method is commonly referred to as two-step data-domain-decomposition (DDD) method. As step (ii) involves the inversion of a straightforward linear system as in conventional CT, step (i) constitutes the key step of the DDD method in which a nonlinear data model needs to be inverted. Although many algorithms can invert the nonlinear model accurately, analysis of its solution existence, uniqueness and stability remains difficult. In recent years, several works studied the solution properties such as [23, 2, 4, 11, 16].
The other is that the scanning geometric parameters are inconsistent or mismatched under different energy spectra. Namely, there is at least an X-ray path that all the X-ray beams under some energy spectrum do not pass through it. We call this case the geometric-inconsistency MSCT. A more extreme example is that for every X-ray path, there is always an energy spectrum under which no X-ray beams pass through it, which can be seen as completely geometric-inconsistency MSCT. The configurations of many commercial devices fall into this situation, such as the fast kV-switching DECT of GE [27].
The development of accurate and fast reconstruction algorithms for this case is more challenging than the geometric-consistency case. The authors in [35] matched the originally mismatched neighboring projections of high- and low-energy spectra in DECT, and then proposed a two-step DDD method to invert those one-view-deviated projection pairs. Similarly, using the matched projections as in [35], an iterative filtered back-projection (FBP) reconstruction method was proposed in [24], which applies Newton’s method with only once iteration and FBP alternately. However, due to the different geometric parameters of the matched projections, these approaches cannot result in accurate image reconstruction. In the last two decades, progress has been made in developing one-step methods for accurate image reconstruction in MSCT. To reconstruct the basis images directly, such kind of method often needs to devise an optimization model from a certain perspective (e.g., statistics) and then develops an iterative scheme to solve that model [13, 6, 33, 25, 5, 7, 28, 8, 15], or considers the problem as solving a large-scale system of nonlinear equations [34, 14]. Because of the nonlinear data model in MSCT, the optimization models designed are generally non-convex. While these methods can often achieve high image quality through a large number of iterations, they are time-consuming, and their convergence analyses were mostly not considered.
In this article, we aim to develop an image reconstruction algorithm of the one-step type that takes into account both accuracy and efficiency, has mathematical guarantee, and is suitable for the geometric-inconsistency MSCT.
The rest of this paper is organized as follows. In section 2, we introduce the forward data models of MSCT. In section 3, we propose a novel algorithm to solve the image reconstruction problem in geometric-inconsistency MSCT. We then prove the convergence of the proposed algorithm in section 4. In section 5, numerical experiments are carried out to illustrate the performance of the proposed algorithm. The flexibility and extensibility of the proposed algorithm are discussed in section 6. Finally, we conclude the paper in section 7.
2 Forward data models in multispectral CT
In this section, the forward data models in MSCT are described briefly, including the continuous-to-continuous (CC)-data model and the discrete-to-discrete (DD)-data model [3, 19, 7, 16].
2.1 CC-data model
The data in MSCT are acquired from an object scanned with multiple X-ray energy spectra $\{s^{[q]}\}_{q=1}^{Q}$, where $s^{[q]}$ is the normalized nonnegative distributed function of the $q$-th energy spectrum, and $Q$ denotes the number of used distinct spectra.
For some given spectrum [^1] The effective spectrum of a CT system is indeed the product of the X-ray-tube spectrum and detector-energy response., data can be detected along the X-ray specified by $l(\theta^{[q]})$, where $\theta^{[q]}$ determines the ray direction. The detected data in the absence of other physical factors can be modeled as
$g_{l(\theta^{[q]})}=\ln\int_{0}^{E_{\text{max}}}s^{[q]}(E)\exp\left(-\int_{l(\theta^{[q]})}\mu(\boldsymbol{y},E)\mathrm{d}l\right)\mathrm{d}E,$ (1)
where $s^{[q]}$ is nonnegative and satisfying
$\int_{0}^{E_{\text{max}}}s^{[q]}(E)\mathrm{d}E=1,$
and $\mu(\boldsymbol{y},E)$ represents the energy-dependent LAC of interest at spatial position $\boldsymbol{y}\in\mathbb{R}^{n}$ ($n=2$ or $3$) and energy $E\in[0,E_{\text{max}}]$, and $\mathbb{R}^{n}$ the real Euclidean space, $E_{\text{max}}$ the maximum energy in the scan. The function $\mu$ of LAC is decomposed often into a linear combination of basis images $\{f_{d}\}_{d=1}^{D}$ as
$\mu(\boldsymbol{y},E)=\sum_{d=1}^{D}b_{d}(E)f_{d}(\boldsymbol{y}),$ (2)
where $b_{d}(E)$ denotes the expansion coefficient, and $D$ the number of basis images. For example, for the typical X-ray energy range in diagnostic DECT, $D=2$ is chosen often as photoelectric effect and Compton scatter are the dominant factors contributing to $\mu$; $b_{d}$ denotes the known function of mass attenuation coefficient (MAC) of the $d$-th basis material (e.g., water and bone); and $f_{d}$ represents the material-equivalent density function of the $d$-th basis material.
Substituting 2 into 1, we obtain
$g_{l(\theta^{[q]})}=\ln\int_{0}^{E_{\text{max}}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)\int_{l(\theta^{[q]})}f_{d}(\boldsymbol{y})\mathrm{d}l\right)\mathrm{d}E.$ (3)
The formula in 3 is referred to as the CC-data model because $\theta^{[q]}$, $E$, and $\boldsymbol{y}$ are continuous variables.
2.2 DD-data model
In practice, data can be detected only for discrete energies and rays, and an image is reconstructed only on an array of discrete pixels/voxels. Hence, it is possible to develop a DD-data model, which may be constructed by implementing a specific discrete scheme into the CC-data model in 3. However, in general, there is no direct connection between the two models.
Assume that the energy domain $[0,E_{\text{max}}]$ is divided into $M$ intervals with equal size $\Delta_{E}$; the data $g_{j}^{[q]}$ are detected for discrete directions $\theta_{j}^{[q]}$ with $1\leq j\leq J_{q}$ under the $q$-th spectrum; each basis image $f_{d}$ is discretized by an array of size $I$. Then a DD-data model is obtained as
$g_{j}^{[q]}=\ln\sum_{m=1}^{M}s^{[q]}_{m}\exp\left(-\sum_{d=1}^{D}b_{dm}\sum_{i=1}^{I}p^{[q]}_{ji}f_{di}\right)\quad\text{for $1\leq j\leq J_{q}$,}$ (4)
where $s^{[q]}_{m}=s^{[q]}(m\Delta_{E})\Delta_{E}$ satisfying $\sum_{m=1}^{M}s^{[q]}_{m}=1$, $b_{dm}=b_{d}(m\Delta_{E})\Delta_{E}$, $f_{di}$ denotes the gray value of the $d$-th discrete basis image at voxel $i$, and $p^{[q]}_{ji}$ is often taken as the intersection length of ray $j$ with pixel/voxel $i$ of the image array [9]. Here the DD-data model 4 can be seen as a specific discrete scheme for the CC-data model 3.
3 The proposed algorithm
In this section, we introduce the MSCT image reconstruction problems and propose new algorithms for the continuous and discrete settings, which correspond to the CC-data model 3 and the DD-data model 4, respectively.
3.1 The continuous setting
Here we give the MSCT image reconstruction problem in continuous setting based on the CC-data model 3, and then present the corresponding algorithm.
3.1.1 Continuous MSCT image reconstruction problem
For simplicity, define $\mathbb{T}^{n}:=\{(\boldsymbol{\omega},\boldsymbol{x})|\boldsymbol{\omega}\in\mathbb{S}^{n-1},\boldsymbol{x}\in\boldsymbol{\omega}^{\bot}\}$, and
$\mathcal{P}f(\boldsymbol{\omega},\boldsymbol{x}):=\int_{-\infty}^{+\infty}f(\boldsymbol{x}+t\boldsymbol{\omega})\mathrm{d}t,\quad(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n},$ (5)
which is the X-ray transform of a certain function $f$ defined on $\mathbb{R}^{n}$, equivalently the Radon transform if $n=2$ [29]. Here $\mathbb{S}^{n-1}$ represents the unit sphere in $\mathbb{R}^{n}$, and $\boldsymbol{\omega}^{\bot}$ denotes the hyperplane perpendicular to $\boldsymbol{\omega}$.
As the X-ray $l(\theta^{[q]})$ in 3 can be parametrized by $(\boldsymbol{\omega},\boldsymbol{x})$, we rewrite 5 as
$\mathcal{P}f(\boldsymbol{\omega},\boldsymbol{x})=\int_{l(\theta^{[q]})}f(\boldsymbol{y})\mathrm{d}l.$ (6)
Using 6, the right-hand side of 3 can be rewritten as
$K^{[q]}(\boldsymbol{f}):=\ln\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)\mathrm{d}E,$ (7)
where $\boldsymbol{f}=[f_{1},\ldots,f_{D}]^{\boldsymbol{\mathsf{T}}}$.
Consequently, the aim of the continuous MSCT image reconstruction is to reconstruct $\boldsymbol{f}$ from the measured data functions $\{g^{[q]}\}_{q=1}^{Q}$ such that
$K^{[q]}(\boldsymbol{f})=g^{[q]}\quad\text{for}~{}q=1,\ldots,Q.$ (8)
Note that 8 can be expressed as the following compact form
$\boldsymbol{K}(\boldsymbol{f})=\boldsymbol{g},$ (9)
where $\boldsymbol{K}(\boldsymbol{f})=[K^{[1]}(\boldsymbol{f}),\ldots,K^{[Q]}(\boldsymbol{f})]^{\boldsymbol{\mathsf{T}}}$ and $\boldsymbol{g}=[g^{[1]},\ldots,g^{[Q]}]^{\boldsymbol{\mathsf{T}}}$.
3.1.2 The proposed algorithm for the continuous problem
Here we propose the algorithm AFIRE for solving the continuous MSCT image reconstruction problem.
Before that, we introduce the required notation. Let $\mathcal{X}$ be a normed vector space with norm $\|\cdot\|_{\mathcal{X}}$, and
$\mathcal{X}^{D}:=\underbrace{\mathcal{X}\times\cdots\times\mathcal{X}}_{D}$
be the Cartesian product of the space $\mathcal{X}$ with the norm defined by
$\|\boldsymbol{u}\|_{{\mathcal{X}}^{D}}:=\left(\sum_{d=1}^{D}\|u_{d}\|^{2}_{\mathcal{X}}\right)^{1/2},$
[…]
Denote the X-ray transform in 5 by the mapping
$\mathcal{P}:\mathcal{F}=H_{0}^{-1/2}(\Omega)\longrightarrow\tilde{\mathcal{G}}=\mathcal{P}(\mathcal{F})\subset\mathcal{L}^{2}(\mathbb{T}^{n})=\mathcal{G}.$
By [30, theorem 2.18], it is known that $\mathcal{P}$ is invertible.
First of all, we have the following proposition.
Proposition 1.
Let $\{s^{[q]}\}_{q=1}^{Q}$ be nonnegative normalized functions and $\{b_{d}\}_{d=1}^{D}$ be nonnegative bounded functions. Then for each $q$, the functional in 7 can be represented as a mapping $K^{[q]}:\mathcal{F}^{D}\rightarrow\mathcal{G}$ and is Gâteaux differentiable, and its Gâteaux derivative at $\boldsymbol{f}\in\mathcal{F}^{D}$ for $\boldsymbol{h}=[h_{1},\ldots,h_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathcal{F}^{D}$ is computed as
$\mathcal{D}K^{[q]}(\boldsymbol{f},\boldsymbol{h})=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}h_{d}\right)\mathrm{d}E,$ (10)
where
$\Phi^{[q]}(\boldsymbol{f};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)\mathrm{d}E}.$ (11)
Moreover, both the $K^{[q]}$ and $\boldsymbol{K}:\mathcal{F}^{D}\rightarrow\mathcal{G}^{Q}$ are Fréchet differentiable.
Proof.
See section A.1 for the proof. ∎
Our insight
We start with analyzing the Gâteaux derivative of $K^{[q]}$ in 10. Notice that the $\Phi^{[q]}(\boldsymbol{f};E)$ in 11 is actually related to the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$ since the term $\mathcal{P}f_{d}$ depends on that parameter as in 6. For a certain real constant set $\{C_{d}\}_{d=1}^{D}$, if there exists one $\hat{\boldsymbol{f}}=[\hat{f}_{1},\ldots,\hat{f}_{D}]^{\boldsymbol{\mathsf{T}}}$ such that
$\mathcal{P}\hat{f}_{d}(\boldsymbol{\omega},\boldsymbol{x})\equiv C_{d},\quad\forall(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n},$ (12)
then $\Phi^{[q]}(\hat{\boldsymbol{f}};E)$ becomes to be independent of the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})$. Together with the linearity of $\mathcal{P}$, the Gâteaux derivative at $\hat{\boldsymbol{f}}$ for $\boldsymbol{h}$ can be rewritten as
$\mathcal{D}K^{[q]}(\hat{\boldsymbol{f}},\boldsymbol{h})=-\mathcal{P}\left(\sum_{d=1}^{D}\varphi^{[q]}_{d}h_{d}\right).$ (13)
Here
$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E,$ (14)
where
$\Phi^{[q]}(\hat{\boldsymbol{f}};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}\right)\mathrm{d}E}.$
Obviously, perhaps the most simple example is taking $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathcal{F}^{D}$ for $C_{d}=0$, $d=1,\ldots,D$ that satisfies the condition of 12. There may be other points that satisfy this condition, but searching for them is beyond the scope of the current research.
Following the insight above, the Gâteaux derivative of $\boldsymbol{K}$ at the $\hat{\boldsymbol{f}}$ determined by 12 can be given as the mapping below
$\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot):\mathcal{F}^{D}$ $\longrightarrow\tilde{\mathcal{G}}^{Q}$
$\boldsymbol{h}$ $\longmapsto-\boldsymbol{\mathcal{P}}(\boldsymbol{\varphi}\boldsymbol{h}),$ (15)
[…]
To proceed, we study the invertibility of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$. We have the following result.
Proposition 2.
Assume that the conditions in proposition 1 hold, and the matrix $\boldsymbol{\varphi}$ is invertible. Then the $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ in section 3.1.2 is invertible.
Proof.
Since $\mathcal{P}$ is invertible from $\mathcal{F}$ to $\tilde{\mathcal{G}}$, $\boldsymbol{\mathcal{P}}$ is invertible from $\mathcal{F}^{D}$ to $\tilde{\mathcal{G}}^{Q}$ when $Q=D$. Combining with the invertibility of $\boldsymbol{\varphi}$, we obtain the invertibility of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$. ∎
Remark 1.
The invertibility of $\boldsymbol{\varphi}$ implies $Q=D$, which means that the number of used energy spectra should be equal to the number of basis images. Actually, this necessary condition is commonly satisfied in clinical applications, for example, $Q=D=2$ in DECT.
Moreover, denote $\mathcal{P}^{-1}$ as the inverse of $\mathcal{P}$, and naturally
$\boldsymbol{\mathcal{P}}^{-1}=\operatorname*{\texttt{diag}}(\mathcal{P}^{-1},\ldots,\mathcal{P}^{-1})$
is the inverse of $\boldsymbol{\mathcal{P}}$. From proposition 2, we can obtain the inverse mapping of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ as
$(\mathcal{D}\boldsymbol{K})^{-1}(\hat{\boldsymbol{f}},\cdot):\tilde{\mathcal{G}}^{Q}$ $\longrightarrow\mathcal{F}^{D}$
$\boldsymbol{r}$ $\longmapsto-\boldsymbol{\varphi}^{-1}\boldsymbol{\mathcal{P}}^{-1}(\boldsymbol{r}).$ (16)
Hence, by leveraging the simplified Newton method [10], the proposed iterative scheme for solving 9 can be formulated as
$\boldsymbol{f}^{k+1}=\boldsymbol{f}^{k}-\boldsymbol{\varphi}^{-1}\boldsymbol{\mathcal{P}}^{-1}\bigl{(}\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k})\bigr{)}.$ (17)
Note that 17 is well-defined if $\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k})\in\tilde{\mathcal{G}}^{Q}$ for $\boldsymbol{f}^{k}\in\mathcal{F}^{D}$. Based on the iterative scheme in 17, we present the algorithm AFIRE for solving the continuous MSCT image reconstruction problem in algorithm 1.
Algorithm 1 The proposed algorithm for solving the continuous problem in 8 (or 9).
1:Initialize: Given the measured data $\{g^{[q]}\}_{q=1}^{Q}$, the energy spectra distributions $\{s^{[q]}\}_{q=1}^{Q}$, the MAC functions $\{b_{d}\}_{d=1}^{D}$ and a certain real constant set $\{C_{d}\}_{d=1}^{D}$. Initialize the basis image $\boldsymbol{f}^{0}$. Let $k\leftarrow 0$.
[…]
3:Loop:
4: Compute the residual
$\begin{bmatrix}\varLambda_{1}^{k}\\ \vdots\\ \varLambda_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}g^{[1]}-K^{[1]}(\boldsymbol{f}^{k})\\ \vdots\\ g^{[Q]}-K^{[Q]}(\boldsymbol{f}^{k})\end{bmatrix}.$
5: Update $\boldsymbol{f}^{k+1}$ by
$\begin{bmatrix}f_{1}^{k+1}\\ \vdots\\ f_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}f_{1}^{k}\\ \vdots\\ f_{D}^{k}\end{bmatrix}-\boldsymbol{\varphi}^{-1}\begin{bmatrix}\mathcal{P}^{-1}\varLambda_{1}^{k}\\ \vdots\\ \mathcal{P}^{-1}\varLambda_{Q}^{k}\end{bmatrix}.$
6: If some given termination condition is satisfied, then output $\boldsymbol{f}^{k+1}$;
otherwise, let $k\leftarrow k+1$, goto Loop.
Remark 2.
In algorithm 1, one can specifically choose the $C_{d}=0$ for $d=1,\ldots,D$ with $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathcal{F}^{D}$ to fulfill the condition in 12. The corresponding formula of 14 becomes
$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}s^{[q]}(E)b_{d}(E)\mathrm{d}E.$
3.2 The discrete setting
Here, we start from the DD-data model 4 and then present the associated algorithm for solving the geometric-inconsistency full-scan MSCT image reconstruction problem.
3.2.1 Discrete MSCT image reconstruction problem
As described in section 2.2, for the discrete case, the forward operator of ray $j$ with the $q$-th energy spectrum may be formulated as
$K_{j}^{[q]}(\mathbf{f})=\ln\sum_{m=1}^{M}s^{[q]}_{m}\exp\left(-\sum_{d=1}^{D}b_{dm}\sum_{i=1}^{I}p^{[q]}_{ji}f_{di}\right).$ (18)
Here $\mathbf{f}:=[\mathbf{f}_{1}^{\boldsymbol{\mathsf{T}}},\ldots,\mathbf{f}_{D}^{\boldsymbol{\mathsf{T}}}]^{\boldsymbol{\mathsf{T}}}$, where $\mathbf{f}_{d}=[f_{d1},\ldots,f_{dI}]^{\boldsymbol{\mathsf{T}}}$ denotes the $d$-th discretized basis image. Hence, the discrete MSCT image reconstruction problem is to determine $\mathbf{f}$ from the measured data $\{g_{j}^{[q]}\}_{j,q=1}^{J_{q},Q}$ by the system of nonlinear equations as below
$K_{j}^{[q]}(\mathbf{f})=g_{j}^{[q]}\quad\text{for}~{}q=1,\ldots,Q~{}\text{and}~{}j=1,\ldots,J_{q}.$ (19)
The compact form of 19 can be expressed as
$\mathbf{K}(\mathbf{f})=\mathbf{g}.$ (20)
Here
$\mathbf{K}(\mathbf{f})=\begin{bmatrix}\mathbf{K}^{[1]}(\mathbf{f})\\ \vdots\\ \mathbf{K}^{[Q]}(\mathbf{f})\end{bmatrix},\mathbf{K}^{[q]}(\mathbf{f})=\begin{bmatrix}K_{1}^{[q]}(\mathbf{f})\\ \vdots\\ K_{J_{q}}^{[q]}(\mathbf{f})\end{bmatrix},\mathbf{g}=\begin{bmatrix}\mathbf{g}^{[1]}\\ \vdots\\ \mathbf{g}^{[Q]}\end{bmatrix},\mathbf{g}^{[q]}=\begin{bmatrix}g^{[q]}_{1}\\ \vdots\\ g^{[q]}_{J_{q}}\end{bmatrix}.$
3.2.2 The proposed algorithm for the discrete problem
For brevity, let us introduce the following notation, for $q=1,\ldots,Q$ and $d=1,\ldots,D$,
$\mathbf{p}_{j}^{[q]}$ $=[p_{j1}^{[q]},\ldots,p_{jI}^{[q]}]^{\boldsymbol{\mathsf{T}}},\quad\mathbf{b}_{d}=[b_{d1},\ldots,b_{dM}]^{\boldsymbol{\mathsf{T}}},\quad\mathbf{s}^{[q]}=[s_{1}^{[q]},\ldots,s_{M}^{[q]}]^{\boldsymbol{\mathsf{T}}},$
$z_{jm}^{[q]}(\mathbf{f})$ $=-\sum_{d=1}^{D}b_{dm}(\mathbf{p}_{j}^{[q]})^{\boldsymbol{\mathsf{T}}}\mathbf{f}_{d},\quad\exp\bigl{(}\mathbf{z}_{j}^{[q]}(\mathbf{f})\bigr{)}=\bigl{[}\exp\bigl{(}z_{j1}^{[q]}(\mathbf{f})\bigr{)},\ldots,\exp\bigl{(}z_{jM}^{[q]}(\mathbf{f})\bigr{)}\bigr{]}^{\boldsymbol{\mathsf{T}}}.$
By 18, the gradient of $K_{j}^{[q]}(\mathbf{f})$ is computed as
$\nabla K_{j}^{[q]}(\mathbf{f})=-\begin{bmatrix}\bigl{(}\mathbf{w}_{j}^{[q]}(\mathbf{f})^{\boldsymbol{\mathsf{T}}}\mathbf{b}_{1}\bigr{)}\mathbf{p}_{j}^{[q]}\\ \vdots\\ \bigl{(}\mathbf{w}_{j}^{[q]}(\mathbf{f})^{\boldsymbol{\mathsf{T}}}\mathbf{b}_{D}\bigr{)}\mathbf{p}_{j}^{[q]}\end{bmatrix},$ (21)
where
$\mathbf{w}_{j}^{[q]}(\mathbf{f})=\frac{\mathbf{s}^{[q]}\odot\exp\bigl{(}\mathbf{z}_{j}^{[q]}(\mathbf{f})\bigr{)}}{(\mathbf{s}^{[q]})^{\boldsymbol{\mathsf{T}}}\exp\bigl{(}\mathbf{z}_{j}^{[q]}(\mathbf{f})\bigr{)}},$ (22)
and $\odot$ denotes the operation of the element-wise product.
Our insight
Similar to the continuous setting, for a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$, if there exists $\hat{\mathbf{f}}=[\hat{\mathbf{f}}_{1}^{\boldsymbol{\mathsf{T}}},\ldots,\hat{\mathbf{f}}_{D}^{\boldsymbol{\mathsf{T}}}]^{\boldsymbol{\mathsf{T}}}$ such that
$\bigl{(}\mathbf{p}_{j}^{[q]}\bigr{)}^{\boldsymbol{\mathsf{T}}}\hat{\mathbf{f}}_{d}\equiv C_{d}^{[q]},\quad\forall j\in\{1,\ldots,J_{q}\},$ (23)
then, using 23, we have
$z_{jm}^{[q]}(\hat{\mathbf{f}})=-\sum_{d=1}^{D}b_{dm}C_{d}^{[q]},$
and hence
$\exp\bigl{(}\mathbf{z}_{j}^{[q]}(\hat{\mathbf{f}})\bigr{)}=\Bigl{[}\exp\Bigl{(}-\sum_{d=1}^{D}b_{d1}C_{d}^{[q]}\Bigr{)},\ldots,\exp\Bigl{(}-\sum_{d=1}^{D}b_{dM}C_{d}^{[q]}\Bigr{)}\Bigr{]}^{\boldsymbol{\mathsf{T}}}.$ (24)
By 24, from 22, it is easy to figure out that
$\mathbf{w}_{j}^{[q]}(\hat{\mathbf{f}})=\frac{\displaystyle\mathbf{s}^{[q]}\odot\exp\bigl{(}-\sum_{d=1}^{D}b_{dm}C_{d}^{[q]}\bigr{)}}{\displaystyle(\mathbf{s}^{[q]})^{\boldsymbol{\mathsf{T}}}\exp\bigl{(}-\sum_{d=1}^{D}b_{dm}C_{d}^{[q]}\bigr{)}},$
which is independent of the index $j$ (namely, the concrete ray). For short, let
$\hat{\mathbf{w}}^{[q]}:=\mathbf{w}_{j}^{[q]}(\hat{\mathbf{f}}).$
Note that probably the most simple example satisfying the condition of 23 is taking $\hat{\mathbf{f}}=\boldsymbol{0}$, which leads to $C_{d}^{[q]}=0$ for any indices $q$ and $d$. But searching for the other points is beyond the scope of this work.
Considering 21 at the $\hat{\mathbf{f}}$ in 23, we obtain that
$\nabla K_{j}^{[q]}(\hat{\mathbf{f}})=-\begin{bmatrix}\phi^{[q]}_{1}\mathbf{p}_{j}^{[q]}\\ \vdots\\ \phi^{[q]}_{D}\mathbf{p}_{j}^{[q]}\end{bmatrix},$
where
$\phi^{[q]}_{d}:=(\hat{\mathbf{w}}^{[q]})^{\boldsymbol{\mathsf{T}}}\mathbf{b}_{d}=\frac{\displaystyle\sum_{m=1}^{M}b_{dm}s_{m}^{[q]}\exp\bigl{(}-\sum_{d=1}^{D}b_{dm}C_{d}^{[q]}\bigr{)}}{\displaystyle\sum_{m=1}^{M}s_{m}^{[q]}\exp\bigl{(}-\sum_{d=1}^{D}b_{dm}C_{d}^{[q]}\bigr{)}}.$ (25)
Then the Jacobian matrix of $\mathbf{K}^{[q]}$ at $\hat{\mathbf{f}}$ can be written as the following form
$\nabla\mathbf{K}^{[q]}(\hat{\mathbf{f}})=-\begin{bmatrix}\phi^{[q]}_{1}\mathbf{P}_{q}&\cdots&\phi^{[q]}_{D}\mathbf{P}_{q}\end{bmatrix},$
where $\mathbf{P}_{q}=[\mathbf{p}_{1}^{[q]},\dots,\mathbf{p}_{J_{q}}^{[q]}]^{\boldsymbol{\mathsf{T}}}$ is the projection matrix under the $q$-th energy spectrum. Furthermore, the associated Jacobian matrix of $\mathbf{K}$ at $\hat{\mathbf{f}}$ is
$\nabla\mathbf{K}(\hat{\mathbf{f}})=-\begin{bmatrix}\phi^{[1]}_{1}\mathbf{P}_{1}&\cdots&\phi^{[1]}_{D}\mathbf{P}_{1}\\ \vdots&\ddots&\vdots\\ \phi^{[Q]}_{1}\mathbf{P}_{Q}&\cdots&\phi^{[Q]}_{D}\mathbf{P}_{Q}\end{bmatrix}=-\mathbf{P}\otimes\boldsymbol{\phi},$ (26)
where $\mathbf{P}=\operatorname*{\texttt{diag}}(\mathbf{P}_{1},\ldots,\mathbf{P}_{Q})$ denotes a diagonal matrix composed of $Q$ projection matrices, $\boldsymbol{\phi}=\bigl{(}\phi^{[q]}_{d}\bigr{)}_{Q\times D}$ represents the constant matrix with entries $\phi^{[q]}_{d}$, and the operation $\otimes$ represents the block multiplication between two matrices.
As the previous continuous situation, we can also study the invertibility of $\nabla\mathbf{K}(\hat{\mathbf{f}})$.
Proposition 3.
Assume that $Q=D$, and $\boldsymbol{\phi}$, $\mathbf{P}_{q}$ for $q=1,\ldots,Q$ are invertible. Then $\nabla\mathbf{K}(\hat{\mathbf{f}})$ is invertible.
Proof.
Since $Q=D$, using the invertibility of matrix, from 26, it is easy to obtain that the inverse of $\nabla\mathbf{K}(\hat{\mathbf{f}})$ is
$\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}^{-1}=-\boldsymbol{\phi}^{-1}\otimes\mathbf{P}^{-1},$
where $\mathbf{P}^{-1}=\operatorname*{\texttt{diag}}(\mathbf{P}_{1}^{-1},\ldots,\mathbf{P}_{Q}^{-1})$. ∎
Again, by leveraging the simplified Newton method as before, the proposed iterative scheme for solving the discrete problem in 19 (or 20) can be formulated as
$\mathbf{f}^{k+1}=\mathbf{f}^{k}-\boldsymbol{\phi}^{-1}\otimes\mathbf{P}^{-1}\bigl{(}\mathbf{g}-\mathbf{K}(\mathbf{f}^{k})\bigr{)}.$ (27)
More clearly, based on 27, we present the proposed algorithm AFIRE for solving 19 (or 20) in algorithm 2, which may also be seen as a specific discretized counterpart of algorithm 1.
Algorithm 2 The proposed algorithm for solving the discrete problem in 19 (or 20).
1:Initialize: Given the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$. Initialized the basis images $\{\mathbf{f}_{d}^{0}\}_{d=1}^{D}$. Let $k\leftarrow 0$.
2:Compute the constant matrix $\boldsymbol{\phi}=\bigl{(}\phi^{[q]}_{d}\bigr{)}_{Q\times D}$ from 25 and the inverse $\boldsymbol{\phi}^{-1}$.
3:Loop:
4: Compute the residual
$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
5: Update $\mathbf{f}^{k+1}$ by
$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{1}^{-1}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{Q}^{-1}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$
6: If some given termination condition is satisfied, then output $\mathbf{f}^{k+1}$;
otherwise, let $k\leftarrow k+1$, goto Loop.
Remark 3.
In algorithm 2, one can particularly choose the $\{C_{d}^{[q]}=0\}_{d,q=1}^{D,Q}$ with $\hat{\mathbf{f}}=\boldsymbol{0}$ to fulfill the condition in 23. Then, the corresponding formula of 25 becomes
$\phi^{[q]}_{d}=\sum_{m=1}^{M}b_{dm}s^{[q]}_{m}.$
As a matter of fact, the inverse of the matrix $\mathbf{P}$ in 27 is difficult to calculate accurately, as all of its diagonal elements $\mathbf{P}_{1},\ldots,\mathbf{P}_{Q}$ may not even be invertible square matrices, even if they are, the computational cost of calculating the inverse of a large-scale matrix is very high. Therefore, we can only obtain an approximation for the inverse of $\mathbf{P}$, and the actual iterative scheme used in algorithm 2 should be
$\mathbf{f}^{k+1}=\mathbf{f}^{k}-\boldsymbol{\phi}^{-1}\otimes\mathbf{P}^{+}\bigl{(}\mathbf{g}-\mathbf{K}(\mathbf{f}^{k})\bigr{)},$ (28)
where $\mathbf{P}^{+}=\operatorname*{\texttt{diag}}(\mathbf{P}_{1}^{+},\ldots,\mathbf{P}_{Q}^{+})$ is the approximation for the inverse of $\mathbf{P}$. Here $\mathbf{P}_{q}^{+}$ denotes the approximation for the inverse of $\mathbf{P}_{q}$. Subsequently, we present the actual algorithm AFIRE used for solving the discrete problem of 19 (or 20) in algorithm 3.
Algorithm 3 The actual algorithm used for solving the discrete problem 19 (or 20).
1:Initialize: Given the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$. Initialized the basis images $\{\mathbf{f}_{d}^{0}\}_{d=1}^{D}$. Let $k\leftarrow 0$.
2:Compute the constant matrix $\boldsymbol{\phi}=\bigl{(}\phi^{[q]}_{d}\bigr{)}_{Q\times D}$ from 25 and the inverse $\boldsymbol{\phi}^{-1}$.
3:Loop:
4: Compute the residual
$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
5: Update $\mathbf{f}^{k+1}$ by
$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{\!1}^{+}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{\!Q}^{+}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$
6: If some given termination condition is satisfied, then output $\mathbf{f}^{k+1}$;
otherwise, let $k\leftarrow k+1$, goto Loop.
Remark 4.
In practice, all the $\mathbf{P}_{q}^{+}$, the approximation for the inverse of $\mathbf{P}_{q}$, $q=1,\dots,Q$, are independent of each other, which can be calculated in parallel by the iterative algorithms, such as algebraic reconstruction technique, conjugate gradient, and the other methods derived from optimization models. Moreover, the full-scan case in this study makes the method of FBP be applicable and efficient.
4 Convergence analysis
In this section, we study the convergence of the proposed algorithms that given in algorithms 1, 2 and 3. Assuming that both the solutions exist to the continuous problem 9 and the discrete one 20, and we refer to $\boldsymbol{f}^{*}$ and $\mathbf{f}^{*}$ as the solutions, respectively.
Let $\|\cdot\|_{F}$ and $\sigma_{\min}(\cdot)$ be the Frobenius norm and the minimum singular value of a matrix, respectively. If $\sigma_{\min}(\cdot)$ is nonvanishing, define the scaled condition number by $\kappa_{F}(\cdot):=\|\cdot\|_{F}/\sigma_{\min}(\cdot)$. Denote $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ as the $1$-norm and Euclidean norm of a vector or the spectral norm of a matrix, respectively.
4.1 Convergence analysis for the continuous setting
To proceed, we give the following lemma.
Lemma 1.
Suppose that the conditions in proposition 1 hold and $\boldsymbol{\varphi}$ is invertible. Assume that there exists a constant $\eta>0$ such that, for any nonnegative normalized function $\rho$ and any indices $q$, $d$,
$\Bigl{|}\int_{0}^{E_{\max}}\left(\rho(E)-\Phi^{[q]}(\hat{\boldsymbol{f}};E)\right)b_{d}(E)\mathrm{d}E\Bigr{|}\leq\eta\Bigl{|}\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E\Bigr{|}.$
Then, for any $\boldsymbol{f},\boldsymbol{h}\in\mathcal{F}^{D}$,
$\|\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}.$ (29)
Proof.
See section A.2 for the proof. ∎
Using the above result, we can establish the global convergence of algorithm 1 and the uniqueness of the solution.
Theorem 1.
Suppose that the conditions in lemma 1 hold and $\eta\kappa_{F}(\boldsymbol{\varphi})<1$, and the iterative scheme 17 is well-defined. Then, for any initial point $\boldsymbol{f}^{0}\in\mathcal{F}^{D}$, the iteration sequence $\{\boldsymbol{f}^{k}\}$ generated by algorithm 1 converges to the unique solution of 9.
Proof.
Using section 3.1.2, the iterative scheme 17 can be rewritten as
$\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k+1}-\boldsymbol{f}^{k})=\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k}).$ (30)
The following estimate is obtained
$\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},$ $\boldsymbol{f}^{k+1}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}$
$=\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k+1}-\boldsymbol{f}^{k})+\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}$
[…]
$=\Big{\|}\int_{0}^{1}\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\mathrm{d}t\Big{\|}_{\mathcal{G}^{D}}$
$\leq\int_{0}^{1}\|\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}\mathrm{d}t$
[…]
With $\eta\kappa_{F}(\boldsymbol{\varphi})<1$, we have
$\big{\|}\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\big{\|}_{\mathcal{G}^{D}}\longrightarrow 0\quad\text{as}~{}k\longrightarrow\infty.$
Based on proposition 2, we know that $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ is invertible. Then, we obtain
$\|\boldsymbol{f}^{k}-\boldsymbol{f}^{*}\|_{\mathcal{F}^{D}}\longrightarrow 0\quad\text{as}~{}k\longrightarrow\infty.$
This completes the proof of the convergence.
Finally, we prove the uniqueness of the solution. If there exists another solution $\bar{\boldsymbol{f}}^{*}$ satisfies 9, according to the above estimate, we have
$\|\boldsymbol{K}(\bar{\boldsymbol{f}}^{*})-\boldsymbol{K}(\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}.$
Thus, $\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}=0$, which implies that $\bar{\boldsymbol{f}}^{*}=\boldsymbol{f}^{*}$. ∎
4.2 Convergence analysis for the discrete setting
For the discrete case, the operator in 20 defined on Euclidean space is obviously continuously differentiable. Similarly, we can establish the following result.
Lemma 2.
Assume that $\boldsymbol{\phi}$ is invertible and $\mathbf{P}$ has full column rank, furthermore, for any $\boldsymbol{\rho}\in\mathbb{R}^{M}_{+}$ in the unit simplex, there exists a constant $\tilde{\eta}>0$ such that for any indices $q$ and $d$,
$\big{|}\mathbf{b}_{d}^{\boldsymbol{\mathsf{T}}}\hat{\mathbf{w}}^{[q]}-\mathbf{b}_{d}^{\boldsymbol{\mathsf{T}}}\boldsymbol{\rho}\big{|}<\tilde{\eta}\big{|}\mathbf{b}_{d}^{\boldsymbol{\mathsf{T}}}\hat{\mathbf{w}}^{[q]}\big{|}.$
Then, for any $\mathbf{f},\mathbf{h}$,
$\|\nabla\mathbf{K}(\hat{\mathbf{f}})\mathbf{h}-\nabla\mathbf{K}(\mathbf{f})\mathbf{h}\|_{2}\leq\tilde{\eta}\kappa_{F}(\mathbf{P}\otimes\boldsymbol{\phi})\|\nabla\mathbf{K}(\hat{\mathbf{f}})\mathbf{h}\|_{2}.$
Proof.
See section A.3 for the proof. ∎
Consider the scheme 28 and 26, and denote the residual in each iteration by
$\boldsymbol{\delta}^{k}=\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{k})-\bigl{(}\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\bigr{)}.$ (31)
For brevity, let
$\gamma=\tilde{\eta}\kappa_{F}(\mathbf{P}\otimes\boldsymbol{\phi}).$
Under proper conditions, we have the following global convergence of algorithm 3.
Theorem 2.
Assume that the conditions in lemma 2 hold and $\gamma<1$. The solution to 20 is unique.
Moreover, for any $k$, $\|\boldsymbol{\delta}^{k}\|_{2}\leq\zeta\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2}$, where $\zeta<(1-\gamma)/(1+\gamma)$. Then, the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 28 converges to the unique solution of 20.
Proof.
By mean value theorem and lemma 2, for any $\mathbf{f}$, we have the following estimate
$\|\mathbf{K}(\mathbf{f})-$ $\mathbf{K}(\mathbf{f}^{*})-\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\|_{2}$
$=\Big{\|}\int_{0}^{1}\nabla\mathbf{K}\bigl{(}t\mathbf{f}+(1-t)\mathbf{f}^{*}\bigr{)}(\mathbf{f}-\mathbf{f}^{*})-\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\mathrm{d}t\Big{\|}_{2}$
$\leq\int_{0}^{1}\|\nabla\mathbf{K}\bigl{(}t\mathbf{f}+(1-t)\mathbf{f}^{*}\bigr{)}(\mathbf{f}-\mathbf{f}^{*})-\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\|_{2}\mathrm{d}t$
$\leq\gamma\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\|_{2}.$ (32)
Since $\gamma<1$ and $\nabla\mathbf{K}(\hat{\mathbf{f}})$ has full column rank, let the $\mathbf{f}$ in section 4.2 be another solution if it exists, the uniqueness of the solution is immediately obtained by section 4.2.
Furthermore, from section 4.2, the triangle inequality implies that
$\|\mathbf{K}(\mathbf{f})-\mathbf{K}(\mathbf{f}^{*})\|_{2}\leq(1+\gamma)\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\|_{2}.$
Hence, by 31 and 4.2, and the above inequality in order, we have
$\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{*})\|_{2}$ $=\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{k})+\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}$
$\leq\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}-\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}+\|\boldsymbol{\delta}^{k}\|_{2}$
$\leq\gamma\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}+\zeta\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2}$
$\leq\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}.$
It follows from $\gamma+\zeta(1+\gamma)<1$ that the sequence $\{\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}\}$ converges to zero. Since $\nabla\mathbf{K}(\hat{\mathbf{f}})$ has full column rank, we obtain the convergence of the iteration sequence. ∎
For the scheme 27, the residual $\boldsymbol{\delta}^{k}$ defined in 31 is vanishing. Following theorem 2, it is easy to obtain the convergence of algorithm 2.
Corollary 1.
Assume that the conditions in lemma 2 hold, $\mathbf{P}$ is invertible and $\gamma<1$. Then the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 27 converges to the unique solution of 20.
Proof.
The proof of theorem 2 can be immediately used here, so we omit the detail. ∎
5 Numerical experiments
We carry out numerical experiments to validate the effectiveness and numerical convergence of the proposed algorithm AFIRE for solving the image reconstruction problem in the geometric-inconsistency DECT in the case of mildly full scan.
The compared algorithms include some existing methods such as the nonconvex primal-dual algorithm (NCPD) [8], the nonlinear Kaczmarz method using cyclic strategy (NKM) [14], and the iterative FBP reconstruction method (IFBP) [24]. Moreover, we consider a practical algorithm that uses linear interpolation firstly for the inconsistent data to obtain the data with consistency, and then employ the two-step DDD method to invert the obtained data, where the method for solving the nonlinear systems in the first step is chosen as Newton’s method. We refer to this algorithm as INTRPL for short.
5.1 Experiment settings
The tests are conducted on a workstation running Python equipped with an Intel i9-13900K @3.0GHz CPU and Nvidia RTX A4000 GPU. The experiment settings for the energy spectra and the MAC are presented as below. We simulate the discrete X-ray energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$ by an open-source X-ray spectrum simulator, SpectrumGUI [31]. Specifically, the used energy spectra pair, as displayed in fig. 1, contains an 80-kV spectrum and a filtered 140-kV spectrum of the GE maxiray 125 X-ray tube, where the filter is a copper of 1-mm width. The MAC of the basis materials, namely, water and bone, are sourced from [20] and used as the decomposition coefficients $\{\mathbf{b}_{d}\}_{d=1}^{D}$. Additionally, we employ two full-scan parallel-beam inconsistent geometries for low and high energy spectra, respectively.
Figure 1: The used energy spectra pair in tests which contains an 80-kV spectrum (red, solid) and a filtered 140-kV spectrum (blue, dashed dot).
With these configurations, we generate the simulated noiseless data by taking the spectra pair, MAC of water and bone, discrete X-ray transform and truth basis images into 18. We utilize the following metrics of relative errors to assess the reconstruction accuracy
$\text{RE}_{\mathbf{f}}^{k}:=\frac{\|\mathbf{f}^{k}-\mathbf{f}^{*}\|_{2}}{\|\mathbf{f}^{*}\|_{2}},\quad\text{RE}_{\mathbf{g}}^{k}:=\frac{\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2}}{\|\mathbf{g}\|_{2}},$
where $\mathbf{f}^{k}$ represents the $k$-th iteration point computed by the used algorithm, $\mathbf{f}^{*}$ the given truth, and $\mathbf{g}$ the simulated noiseless/noisy data. Moreover, we use the following error metrics between adjacent iterations to further evaluate the numerical convergence of the proposed algorithm
$\Delta_{\mathbf{f}}^{k}:=\frac{\|\mathbf{f}^{k+1}-\mathbf{f}^{k}\|_{2}}{\|\mathbf{f}^{k}\|_{2}},\quad\Delta_{\mathbf{g}}^{k}:=\frac{\|\mathbf{K}(\mathbf{f}^{k+1})-\mathbf{K}(\mathbf{f}^{k})\|_{2}}{\|\mathbf{g}\|_{2}}.$
In algorithm 3 of AFIRE, as stated in remarks 3 and 4, we take $\hat{\mathbf{f}}=\boldsymbol{0}$, and implement the $\mathbf{P}_{q}^{+}$, $q=1,\ldots,Q$ by FBP method where the 1-bandwidth Ram–Lak filter is used. Moreover, INTRPL and IFBP also use the same FBP method, and NCPD is used to solve a total variation (TV)-regularized nonlinear least square problem with nonnegative constraint. For NCPD, the regularization parameter is selected as 00 and $10^{-6}$ for the noiseless and noisy data, respectively. Without loss of generality, the initial point for all tested algorithms are set to the zero vector.
5.2 Simulated noiseless data for the forbild phantom
In this test, a forbild phantom is composed of the truth water and bone basis images which are both discretized as $128\!\times\!128$ pixels on the square area $[-5,5]\!\times\![-5,5]$ $\text{cm}^{2}$ and take gray values over $[0,1]$, as shown in row 1 of fig. 5.
The X-ray paths under the different spectra displayed in fig. 1 are completely inconsistent. Specifically, for the 80-kV spectrum, we utilize 384 parallel projections uniformly sampled on $[-7.05,7.05]$ cm for each of the 384 views, where these views are uniformly distributed over the interval $[0,\pi)$ and can be seen as a mildly full scan. Under the filtered 140-kV spectrum, the scanning also has the same number of projections and views, but the views are uniformly distributed over the interval $[\pi/768,\pi/768+\pi)$.
After 100 iterations, we plot the curves of metrics $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in fig. 2.
[…]
It is shown by the metrics of relative error that the proposed algorithm AFIRE has reached numerical convergence and reconstructed high-accuracy result in less than 50 iterations, thereby demonstrating its effectiveness and accuracy for inverting noiseless data. The results also confirm the convergence theory established in theorem 2. In contrast, NCPD and NKM are still far from convergence, and INTRPL can only produce low-accuracy result. Particularly, it is observed that IFBP does not converge.
The metrics of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ over time are also presented in fig. 3. We observe that the proposed algorithm AFIRE exhibits superior performance, demonstrating its advantages in both efficiency and accuracy over the other compared algorithms.
[…]
We further plot the curves of metrics $\Delta_{\mathbf{f}}^{k}$ and $\Delta_{\mathbf{g}}^{k}$ in fig. 4, which also demonstrate the numerical convergence of the proposed algorithm AFIRE.
Figure 4: Metrics $\Delta_{\mathbf{f}}^{k}$ (left) and $\Delta_{\mathbf{g}}^{k}$ (right) are plotted in semi-log scale as functions of iteration numbers for reconstructing basis images of the forbild phantom by the proposed algorithm AFIRE from the noiseless data.
Furthermore, as shown in fig. 5, the reconstructed basis images and the obtained VMIs after 100 iterations are displayed for all the above algorithms except for IFBP, which reveal that the reconstructed results by the proposed algorithm AFIRE are closest to the truths.
Figure 5: From left to right: basis images of water and bone, VMIs at energies 60 keV and 100 keV of the forbild phantom. From top to bottom: the truths, the results after 100 iterations using the proposed algorithm AFIRE, and NCPD, NKM, INTRPL.
5.3 Simulated noisy data for the realistic torso image
For this test, a real patient torso image is decomposed into the truth water and bone basis images which are both discretized as $256\!\times\!256$ pixels on the square area $[-5,5]\!\times\![-5,5]$ $\text{cm}^{2}$ and take gray values over $[0,1]$, as shown in row 1 of fig. 9.
Similarly, the X-ray paths under the different spectra shown in fig. 1 are also completely inconsistent. More precisely, under the 80-kV spectrum, 768 parallel-beam projections uniformly distributed on $[-7.05,7.05]$ cm are taken for each of 768 views, where these views are evenly sampled on the interval $[0,\pi)$. As to the filtered 140-kV spectrum, the same numbers of projections and views are utilized, but the views are uniformly distributed over the interval $[\pi/1536,\pi/1536+\pi)$. Moreover, the 34.3 dB noisy data is generated by adding Gaussian noise onto the noiseless data.
We plot the curves of metrics $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in fig. 6 within 100 iterations. It is obvious that the metrics of relative errors for the proposed algorithm AFIRE exhibit numerical convergence and relatively high accuracy in less than $15$ iterations, demonstrating its effectiveness and stability for addressing noisy data. On the contrary, NCPD and NKM have not converged yet; INTRPL can only produce low-accuracy result; IFBP is divergent after a couple of iterations.
[…]
The metrics of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ over time are presented in fig. 7. It is observed that the proposed algorithm AFIRE exhibits superior performance, demonstrating its advantages in both efficiency and accuracy over the other compared algorithms.
[…]
We also present the curves of metrics $\Delta_{\mathbf{f}}^{k}$ and $\Delta_{\mathbf{g}}^{k}$ in fig. 8, which definitely validate the numerical convergence of the proposed algorithm AFIRE.
Figure 8: Metrics $\Delta_{\mathbf{f}}^{k}$ (left) and $\Delta_{\mathbf{g}}^{k}$ (right) are plotted in semi-log scale as functions of iteration numbers for reconstructing basis images of the real torso image by the proposed algorithm AFIRE from the noisy data.
Moreover, after 100 iterations, the reconstructed basis images and the obtained VMIs are displayed in fig. 9 for all the above algorithms except for IFBP, which reveal that the reconstructed images by the proposed algorithm AFIRE are closest to the truths.
Figure 9: From left to right: basis images of water and bone, VMIs at energies 60 keV and 100 keV of the realistic torso image. From top to bottom: the truths, the results after 100 iterations using the proposed algorithm AFIRE, and NCPD, NKM, INTRPL.
6 Discussion
In this section, we discuss more details about the proposed algorithm AFIRE. Actually, this algorithm can be seen as a framework for solving the image reconstruction problem in MSCT, since there are two main adjustable modules, for instance, including the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$ in algorithm 3 of AFIRE. On the other hand, although this algorithm was originally designed to handle the reconstruction problem of geometric inconsistency between low and high X-ray energy spectra, it can be directly used to solve the problem in geometric-consistency situation. Hence, the flexibility and extensibility may enhance the applicability of the proposed algorithm.
To begin with, we introduce some basic experiment settings. The same energy spectra and MAC are applied as those described in section 5. Differently, the tested water and bone basis images are originated from the realistic CT images which are randomly selected from the dataset of LoDoPaB-CT [22], as shown in row 1 of figs. 11, 12 and 14. All the basis images are discretized as $362\!\times\!362$ pixels on the square area $[-15,15]\!\times\![-15,15]$ $\text{cm}^{2}$.
In the discussion of the two adjustable modules, under the 80-kV spectrum, we employ 1086 parallel projections evenly sampled on $[-21.15,21.15]$ cm for each of the 900 views, where these views are uniformly distributed over the interval $[0,\pi)$. Under the filtered 140-kV spectrum, the scanning uses the same numbers of projections and views, but the views are uniformly distributed over the interval $[\pi/1800,\pi/1800+\pi)$. With these configurations, we generate simulated noiseless data as before and apply algorithm 3 with selecting different modules to solve the corresponding reconstruction problems.
6.1 Selection of the special point $\hat{\mathbf{f}}$
Based on our insights in section 3, in the discrete (continuous) case, obviously, for the real constant set $\{C_{d}^{[q]}=0\}_{d,q=1}^{D,Q}$ ($\{C_{d}=0\}_{d=1}^{D}$), there exists $\hat{\mathbf{f}}=\boldsymbol{0}$ ($\hat{\boldsymbol{f}}=\boldsymbol{0}$) fulfilling the condition 23 (the condition 12). As we can imagine, those are almost the most simple examples satisfying the required conditions.
The results in section 5 have demonstrated the effectiveness and numerical convergence of the proposed algorithm when we set each $C_{d}^{[q]}=0$ with $\hat{\mathbf{f}}=\boldsymbol{0}$. Besides that, we further investigate alternative choices. Notably, the implementation of algorithm 3 depends only on the values of certain constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$, even if it is unknown whether there exists an $\hat{\mathbf{f}}$ to satisfy the condition 23.
To discuss these alternatives, we just generate three random arrays over [0,1] as different selections of $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$ using numpy.random with different seeds, but ignore searching for the concrete special point $\hat{\mathbf{f}}$. Additionally, all the $\mathbf{P}_{q}^{+}$, $q=1,\ldots,Q$, are implemented by the FBP with 1-bandwidth Ram–Lak filter.
As shown in fig. 10, we plot the curves of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in $100$ iterations.
[…]
It is discovered that the convergence behaviors of the proposed algorithm using these three different arrays are almost the same as that using the vanishing one. The basis images and VMIs obtained by the proposed algorithm using these arrays are displayed in fig. 11, and they are visually comparable and are all quite close to the truth images.
Figure 11: From left to right: basis images of water and bone, VMIs at energies 60 keV and 100 keV of a clinical image. From top to bottom: the truths, the results after 100 iterations by the proposed algorithm AFIRE with different selections of $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$.
These results indicate that the convergence, accuracy and efficiency of the proposed algorithm AFIRE remain robust across different choices of $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$, and naturally the selection of the special point $\hat{\mathbf{f}}$ is also flexible. This enhances the applicability of the algorithm for various practical scenarios without compromising performance.
6.2 Implementation of the approximated inverse $\mathbf{P}^{+}$
In section 5, we employed the FBP with 1-bandwidth Ram–Lak filter to implement the approximated inverse $\mathbf{P}^{+}$. Besides that, the proposed algorithm AFIRE also allows us to adopt different implementations for $\mathbf{P}^{+}$. Specifically, we further consider the other two different implementation methods: conjugate gradient method with 20 inner iterations (CG) and limited-memory BFGS method with 60 inner iterations (L-BFGS). Note that, the numbers of the inner iterations are not fixed, which can be modified on demand. All the implementations are running with Operator Discretization Library (ODL) [1].
Here, the numerical experiments are carried out to evaluate the feasibility and impact of the different implementations for $\mathbf{P}^{+}$ on the performance of the proposed algorithm.
We still set $\hat{\mathbf{f}}=\boldsymbol{0}$. After running 100 iterations with algorithm 3, the obtained basis images and VMIs using those different implementations are shown in fig. 12, which are quite close to the corresponding truths.
Figure 12: From left to right: basis images of water and bone, VMIs at energies 60 keV and 100 keV of a clinical image. From top to bottom: the truths, the results after 100 iterations by the proposed algorithm AFIRE with the implementations of FBP, CG and L-BFGS for the approximated inverse $\mathbf{P}^{+}$.
Moreover, as indicated in fig. 13, the metric curves of $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ also reveal the similar convergence behavior among those implementations.
[…]
The implementation that requires inner iterations such as CG and L-BFGS achieves slightly lower error of $\text{RE}_{\mathbf{f}}^{k}$ than the used FBP. On the contrary, as evidenced by the comparison of computing time in table 1, FBP is faster than the other two implementations in single (outer) iteration and when the errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ reach $10^{-5}$. As proved in theorem 2, more precise implementation of $\mathbf{P}^{+}$ may improve convergence rate of the entire algorithm. However, it may increase the computational cost and reduce the efficiency. Therefore, there is a trade-off between the reconstructed accuracy and computational efficiency.
Table 1: The average time (in seconds) of 10 independent runs by the proposed algorithm AFIRE with different implementations for $\mathbf{P}^{+}$ to accomplish various tasks in reconstructing basis images of a clinical image in fig. 12.
  Time (s) FBP CG L-BFGS
  Single (outer) iteration $0.47$ 0.80 3.22
$\text{RE}_{\mathbf{f}}^{k}$ reaches $10^{-5}$ $13.99$ 20.70 86.81
$\text{RE}_{\mathbf{g}}^{k}$ reaches $10^{-5}$ $8.86$ 15.13 61.09
6.3 Extension to the geometric-consistency situation
From the derivation in section 3, we know that the proposed algorithm AFIRE can be naturally extended to the geometric-consistency case. To illustrate this, what is different from before is that the identical equidistant fan-beam geometry is used for both 80-kV and 140-kV spectra. There are totally 900 views uniformly distributed over $[0,2\pi)$ and 1086 equal-width detectors positioned on the opposing curved detector bank.
The proposed algorithm is used to reconstruct the basis images from the noiseless data. In the geometric-consistency case, the two-step DDD method can be used as a benchmark algorithm for comparison. In the first step, to estimate the basis sinograms, we use Newton’s method to solve the associated nonlinear systems in data domain, where the iteration number is 10 and the relative error achieves $10^{-14}$. In the second step, to reconstruct the basis images, the FBP with 1-bandwidth Ram–Lak filter is also applied to invert the obtained basis sinograms.
After running algorithm 3 for 100 iterations with setting $\hat{\mathbf{f}}=\boldsymbol{0}$, the obtained basis images and VMIs are shown in fig. 14, as well as the corresponding results gained by the two-step DDD method. Furthermore, the relative error curves are presented in fig. 15. These comparisons demonstrate that the performance of the proposed algorithm far exceeds the benchmark two-step DDD method. Hence, the proposed algorithm cannot only be successfully extended to deal with the geometric-consistency case of MSCT image reconstruction, but also achieve very high accuracy and efficiency.
Figure 14: From left to right: basis images of water and bone, VMIs at energies 60 keV and 100 keV of a clinical image. From top to bottom: the truths, the results by the proposed algorithm AFIRE after 100 iterations and the benchmark two-step DDD method.
[…]
7 Conclusion
In this work, the image reconstruction algorithm was investigated for the nonlinear MSCT when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched.
We studied the associated CC (DD)-data model, and discovered that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix. Based on the insights, we proposed an accurate and fast image reconstruction algorithm called AFIRE respectively from the continuous, discrete and actual-use perspectives by leveraging the simplified Newton method. Under proper conditions, we further proved the convergence of the proposed algorithm AFIRE.
Furthermore, several numerical experiments were carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in geometric-inconsistency DECT for noiseless and noisy projection data in the case of mildly full scan. Through numerical comparisons, it was shown that the proposed algorithm significantly outperforms some existing methods such as NCPD, NKM, IFBP and INTRPL with respect to the accuracy and efficiency.
We also discussed the flexibility and extensibility of the proposed algorithm. It was demonstrated that the proposed algorithm is robust and flexible with regard to the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$. Moreover, for the geometric-consistency situation, it was shown that the proposed algorithm has significant advantages over the benchmark two-step DDD method.
To the best of our knowledge, when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched in MSCT, this is the first literature that uses the characteristics of the derivative operator (gradient) of the associated nonlinear mapping to construct the algorithm for accurate and fast image reconstruction. Moreover, this methodology may also provide insights into designing the algorithm of the same type for the corresponding low-dose or sparse-view scenarios.
Appendix A Appendix
A.1 The proof of proposition 1
Proof.
For any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$, $K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})$ can be rewritten as the function $H_{q}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ which is given by
$\boldsymbol{z}$ $\longmapsto\ln\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)\mathrm{d}E,$
where $\boldsymbol{z}=[z_{1},\ldots,z_{D}]^{\boldsymbol{\mathsf{T}}}$ and $z_{d}=\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})$. It is easy to compute that
$\frac{\partial H_{q}(\boldsymbol{z})}{\partial z_{d}}=-\int_{0}^{E_{\max}}\frac{s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)}{\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)\mathrm{d}E}b_{d}(E)\mathrm{d}E.$ (33)
By 33 and the nonnegative bounded property of functions $\{b_{d}\}_{d=1}^{D}$, we have
$\|\nabla H_{q}(\boldsymbol{z})\|_{1}=\sum_{d=1}^{D}\left|\frac{\partial H_{q}(\boldsymbol{z})}{\partial z_{d}}\right|\leq\sum_{d=1}^{D}\sup_{E}b_{d}(E).$ (34)
Notice that $K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})=0$ for any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$, and using 34, we obtain
$\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\left|H_{q}(\boldsymbol{z})-H_{q}(0)\right|\\ \leq\left(\sup_{\boldsymbol{z}}\|\nabla H_{q}(\boldsymbol{z})\|_{1}\right)\|\boldsymbol{z}\|_{1}\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|.$ (35)
Consequently, for any $\boldsymbol{f}\in\mathcal{F}^{D}$, by 35, we have
$\|K^{[q]}(\boldsymbol{f})\|_{\mathcal{G}}$ $=\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\Bigl{(}\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|\Bigr{)}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\|\mathcal{P}f_{d}\|_{\mathcal{G}},$
where $\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}$ denotes the Euclidean measure on the hyperplane $\boldsymbol{\omega}^{\perp}$. Since $\mathcal{P}f_{d}\in\mathcal{G}$ for each $d$, we obtain $K^{[q]}(\boldsymbol{f})\in\mathcal{G}$.
For any $\boldsymbol{h}=[h_{1},\ldots,h_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathcal{F}^{D}$, by the chain rule, we compute
$\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}h_{d}\right)\mathrm{d}E,$ (36)
where $\Phi^{[q]}(\boldsymbol{f};E)$ is given as in 11 such that
$\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathrm{d}E=1\quad\text{for}~{}q=1,\ldots,Q.$
Then, from 36, we have the following estimate
$\big{\|}\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}\big{\|}_{\mathcal{G}}$ $=\big{\|}\sum_{d=1}^{D}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)b_{d}(E)\mathcal{P}h_{d}\mathrm{d}E\big{\|}_{\mathcal{G}}$ (37)
$\leq\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\big{\|}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathcal{P}h_{d}\mathrm{d}E\big{\|}_{\mathcal{G}}$
$=\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\left\|\mathcal{P}h_{d}\right\|_{\mathcal{G}}$
$\leq\left(\sup_{d,E}b_{d}(E)\right)C\sum_{d=1}^{D}\|h_{d}\|_{\mathcal{F}}.$
Note that the last inequality follows readily from the result in [30, theorem 2.10], and the $C$ is a positive constant depending only on the $\alpha=-1/2$ and the domain $\Omega$. Hence, using 36 and 37, we conclude that $K^{[q]}$ is Gâteaux differentiable, and its Gâteaux derivative is computed as 10.
Furthermore, from the continuity of $\mathcal{P}$, it is easy to obtain that the mapping $\boldsymbol{f}\mapsto\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$ is continuous. By [12, proposition 3.2.15], $K^{[q]}$ is Fréchet differentiable and its Fréchet derivative is equal to $\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$. By the definition of the norm of Cartesian space, we can further prove that $\boldsymbol{K}$ is also Fréchet differentiable. ∎
A.2 The proof of lemma 1
Proof.
For any $\boldsymbol{f}$ and $(\boldsymbol{\omega},\boldsymbol{x})$, we can define the matrix
$\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x}):=\bigl{(}\psi^{[q]}_{d}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\bigr{)}\in\mathbb{R}^{Q\times D},$
where
$\psi^{[q]}_{d}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})=\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})b_{d}(E)\mathrm{d}E.$
Since $\Phi^{[q]}(\boldsymbol{f};\,\cdot\,)$ as given in 11 is a nonnegative normalized function with respect to $E$, together with the given conditions, we have
$\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\|_{F}^{2}$ $=\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\big{(}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})-\Phi^{[q]}(\hat{\boldsymbol{f}};E)\big{)}b_{d}(E)\mathrm{d}E\right)^{2}$
$\leq\eta^{2}\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E\right)^{2}=\eta^{2}\|\boldsymbol{\varphi}\|_{F}^{2}.$
Subsequently, for any $\boldsymbol{z}\in\mathbb{R}^{D}$, by Cauchy–Schwarz inequality and the property of matrix norm, we have
$\|(\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi})\boldsymbol{z}\|_{2}$ $\leq\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\|_{2}\|\boldsymbol{z}\|_{2}\leq\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\|_{F}\|\boldsymbol{z}\|_{2}$
$\leq\eta\|\boldsymbol{\varphi}\|_{F}\|\boldsymbol{z}\|_{2}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\boldsymbol{\varphi}\boldsymbol{z}\|_{2}.$
Hence, for any $\boldsymbol{f},\boldsymbol{h}\in\mathcal{F}^{D}$, based on the formula of Fréchet derivative of $\boldsymbol{K}$ and the above estimation, we obtain
$\big{\|}\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\big{\|}_{\mathcal{G}^{Q}}^{2}$ $=\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\left\|\left(\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\right)\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})\right\|_{2}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}$
$\leq\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}(\eta\kappa_{F}(\boldsymbol{\varphi}))^{2}\|\boldsymbol{\varphi}\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})\|_{2}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}$
$=(\eta\kappa_{F}(\boldsymbol{\varphi}))^{2}\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}^{2},$
where $\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})=[\mathcal{P}h_{1}(\boldsymbol{\omega},\boldsymbol{x}),\ldots,\mathcal{P}h_{D}(\boldsymbol{\omega},\boldsymbol{x})]^{\boldsymbol{\mathsf{T}}}\in\mathbb{R}^{D}$. The desired result is obtained. ∎
A.3 The proof of lemma 2
Proof.
Notice that for any $\mathbf{f}$, $\mathbf{w}_{j}^{[q]}(\mathbf{f})\geq\boldsymbol{0}$ and $\bigl{(}{\mathbf{w}_{j}^{[q]}(\mathbf{f})}\bigr{)}^{\boldsymbol{\mathsf{T}}}\boldsymbol{1}=1$, which means that $\mathbf{w}_{j}^{[q]}(\mathbf{f})\in\mathbb{R}^{M}_{+}$ is in the unit simplex. Hence, by the gradient calculation in section 3.2.2, and using the given conditions, we obtain
$\big{\|}\nabla\mathbf{K}(\hat{\mathbf{f}})-\nabla\mathbf{K}(\mathbf{f})\big{\|}_{F}^{2}=$ $\sum_{q=1}^{Q}\sum_{j=1}^{J_{q}}\|\nabla K_{j}^{[q]}(\hat{\mathbf{f}})-\nabla K_{j}^{[q]}(\mathbf{f})\|_{2}^{2}$
$=$ $\sum_{q=1}^{Q}\sum_{j=1}^{J_{q}}\sum_{d=1}^{D}|\big{(}\mathbf{w}_{j}^{[q]}(\hat{\mathbf{f}})-\mathbf{w}_{j}^{[q]}(\mathbf{f})\big{)}^{\boldsymbol{\mathsf{T}}}\mathbf{b}_{d}\mathbf{p}^{[q]}_{j}|^{2}$
$=$ $\sum_{q=1}^{Q}\sum_{j=1}^{J_{q}}\sum_{d=1}^{D}|\mathbf{b}_{d}^{\boldsymbol{\mathsf{T}}}{\hat{\mathbf{w}}}^{[q]}-\mathbf{b}_{d}^{\boldsymbol{\mathsf{T}}}{\mathbf{w}_{j}^{[q]}(\mathbf{f})}|^{2}\|\mathbf{p}^{[q]}_{j}\|_{2}^{2}$
$\leq$ $\sum_{q=1}^{Q}\sum_{j=1}^{J_{q}}\sum_{d=1}^{D}\tilde{\eta}^{2}|\mathbf{b}_{d}^{\boldsymbol{\mathsf{T}}}\hat{\mathbf{w}}^{[q]}|^{2}\|\mathbf{p}^{[q]}_{j}\|_{2}^{2}$
$=$ $\tilde{\eta}^{2}\|\nabla\mathbf{K}(\hat{\mathbf{f}})\|_{F}^{2}.$
Consequently, by the property of the matrix norm, the above estimate in order, we have
$\big{\|}\nabla\mathbf{K}(\hat{\mathbf{f}})\boldsymbol{h}-\nabla\mathbf{K}(\mathbf{f})\mathbf{h}\big{\|}_{2}$ $\leq\|\nabla\mathbf{K}(\hat{\mathbf{f}})-\nabla\mathbf{K}(\mathbf{f})\|_{F}\|\mathbf{h}\|_{2}\leq\tilde{\eta}\|\nabla\mathbf{K}(\hat{\mathbf{f}})\|_{F}\|\mathbf{h}\|_{2}$
$\leq\tilde{\eta}\|\nabla\mathbf{K}(\hat{\mathbf{f}})\|_{F}\sigma_{\min}^{-1}\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})\mathbf{h}\|_{2}$
$=\tilde{\eta}\kappa_{F}(\mathbf{P}\otimes\boldsymbol{\phi})\|\nabla\mathbf{K}(\hat{\mathbf{f}})\mathbf{h}\|_{2},$
which completes the proof. ∎

### v2

@@
-AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistency Multispectral CT
+AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistent Multispectral CT
 Abstract
 For nonlinear multispectral computed tomography (CT), accurate and fast image reconstruction is challenging when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched.
-Motivated by this, we propose an accurate and fast algorithm named AFIRE to address such problem in the case of mildly full scan.
-We discover that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix.
-Based on the insights, the AFIRE is proposed respectively from the continuous, discrete and actual-use perspectives by leveraging the simplified Newton method.
+Motivated by this, we propose an Accurate and Fast Image REconstruction (AFIRE) algorithm to address such problems in the case of mildly full scan.
+From the continuous (resp.
+discrete) setting, we discover that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix.
+Based on these insights, the AFIRE algorithm is proposed by leveraging the simplified Newton method.
 Under proper conditions, we establish the convergence theory of the proposed algorithm.
-Furthermore, numerical experiments are also carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in completely geometric-inconsistency dual-energy CT with noiseless and noisy projection data.
+Furthermore, numerical experiments are also carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in completely geometric-inconsistent dual-energy CT with noiseless and noisy projection data.
 Particularly, the proposed algorithm significantly outperforms some state-of-the-art methods in terms of accuracy and efficiency.
@@
 keywords:
-Multispectral CT, dual-energy CT, geometric inconsistency, large-scale nonlinear inverse problem, accurate and fast image reconstruction algorithm, one-step method, convergence analysis
+Multispectral CT, dual-energy CT, geometric inconsistency, accurate and fast image reconstruction algorithm, large-scale nonlinear inverse problem, one-step method, convergence analysis
 1 Introduction
-Basically, multispectral computed tomography (MSCT) is an imaging modality aiming to reconstruct material-specific (basis) images from multiple sets of attenuation data collected with different X-ray energy spectra, which can subsequently be used to form virtual monochromatic images (VMIs) or the spatial distributions of linear attenuation coefficient (LAC) at energies of interest by linear combination [18, 3, 27].
+Multispectral CT (MSCT) is an imaging modality aiming to reconstruct material-specific (basis) images from multiple sets of attenuation data collected with different X-ray energy spectra.
+These basis images can subsequently be used to form virtual monochromatic images (VMIs) or the spatial distributions of linear attenuation coefficient (LAC) at energies of interest by linear combination [20, 4, 29].
 MSCT has attracted increasing attention and research interest due to the development of hardware and the requirement of applications, as well as its advantages in considering the X-ray energy-dependent information over the conventional CT such as artifact suppression, contrast improvement and object separation [29].
@@
 This assumption makes the model linear, which causes the image reconstruction to become a linear inverse problem, but leading to difficulties in artifact suppression and limitations in material differentiation.
-In real-world CT, however, this assumption is not realistic, since the output X-ray beams are polychromatic, which consist of photons with a wide range of energies that distributed as the energy spectrum.
+In real-world CT, however, this assumption is not realistic because the output X-ray beams are polychromatic, consisting of photons with a wide range of energies.
+The distribution of photon energies is referred to as the energy spectrum.
 The LAC of the scanned material depends not only on its own elemental composition but also on the energy of the probing X-ray photons.
@@
 Although development of algorithms for reconstructing basis images in MSCT constitutes an important topic of active research, accurate and fast image reconstruction remains a challenging task, largely because its forward data model is nonlinear.
-In this work, we understand different reconstruction algorithms from the perspective of whether the scanning geometric parameters are consistent or matched under different X-ray energy spectra.
-One situation is that the scanning geometric parameters are consistent under different energy spectra.
-That is to say, for each X-ray path, under every energy spectrum, there is always an X-ray beam passing through it.
-Here this case is called geometric-consistency MSCT, in which the image reconstruction can be accurately realized by a two-step method: step (i) estimates the basis sinograms from measurements through solving a lot of very small-scale independent systems of nonlinear equations, whereas step (ii) reconstructs basis images from the basis sinograms estimated.
+In this work, we understand different reconstruction algorithms from the perspective of scanning geometric parameters under different X-ray energy spectra.
+One situation is that the scanning geometric parameters are consistent or matched under different energy spectra.
+That is to say, for each X-ray path, under every energy spectrum, there is always one X-ray beam passing through it.
+Here this case is called geometric-consistent MSCT, in which the image reconstruction can be accurately realized by a two-step method: step (i) estimates the basis sinograms from measurements through solving a lot of very small-scale independent systems of nonlinear equations along different X-ray paths, for instance, using Newton’s method, whereas step (ii) reconstructs basis images from the basis sinograms estimated, such as using the filtered back-projection (FBP) method.
 This method is commonly referred to as two-step data-domain-decomposition (DDD) method.
-As step (ii) involves the inversion of a straightforward linear system as in conventional CT, step (i) constitutes the key step of the DDD method in which a nonlinear data model needs to be inverted.
+As step (ii) involves the inversion of a linear system as in conventional CT, step (i) constitutes the key step of the DDD method in which a nonlinear data model needs to be inverted.
 Although many algorithms can invert the nonlinear model accurately, analysis of its solution existence, uniqueness and stability remains difficult.
@@
 The other is that the scanning geometric parameters are inconsistent or mismatched under different energy spectra.
-Namely, there is at least an X-ray path that all the X-ray beams under some energy spectrum do not pass through it.
-We call this case the geometric-inconsistency MSCT.
-A more extreme example is that for every X-ray path, there is always an energy spectrum under which no X-ray beams pass through it, which can be seen as completely geometric-inconsistency MSCT.
+Namely, there is at least one X-ray path that all the X-ray beams under some energy spectrum do not pass through it.
+We call this case the geometric-inconsistent MSCT.
+A more extreme example is that for every X-ray path, there is always an energy spectrum under which no X-ray beams pass through it, which can be seen as completely geometric-inconsistent MSCT.
 The configurations of many commercial devices fall into this situation, such as the fast kV-switching DECT of GE [29].
-The development of accurate and fast reconstruction algorithms for this case is more challenging than the geometric-consistency case.
+The development of accurate and fast reconstruction algorithms for this case is more challenging than the geometric-consistent case.
 The authors in [38] matched the originally mismatched neighboring projections of high- and low-energy spectra in DECT, and then proposed a two-step DDD method to invert those one-view-deviated projection pairs.
-Similarly, using the matched projections as in [35], an iterative filtered back-projection (FBP) reconstruction method was proposed in [24], which applies Newton’s method with only once iteration and FBP alternately.
-However, due to the different geometric parameters of the matched projections, these approaches cannot result in accurate image reconstruction.
-In the last two decades, progress has been made in developing one-step methods for accurate image reconstruction in MSCT.
-To reconstruct the basis images directly, such kind of method often needs to devise an optimization model from a certain perspective (e.g., statistics) and then develops an iterative scheme to solve that model [13, 6, 33, 25, 5, 7, 28, 8, 15], or considers the problem as solving a large-scale system of nonlinear equations [34, 14].
-Because of the nonlinear data model in MSCT, the optimization models designed are generally non-convex.
+Similarly, using the matched projections as in [38], an iterative FBP (IFBP) reconstruction method was proposed in [26], which applies a single Newton iteration and FBP alternately.
+However, due to the different geometric parameters of the matched projections, these approaches cannot result in an accurate image reconstruction.
+In the last two decades, progress has been made in developing one-step methods for the image reconstruction in MSCT.
+To reconstruct the basis images directly, such kind of method often needs to devise an optimization model from a theoretical perspective, such as Bayesian inference and variational regularization, and then develop an iterative scheme to solve that model.
+For instance, the authors in [15, 8, 36, 27, 30] employed the maximum a posteriori estimation within Bayesian inference to formulate the optimization models, and the authors in [6, 9, 10, 17] constructed the optimization models from the perspective of variational regularization.
+Because of the nonlinear data model in MSCT, the optimization models designed are generally nonconvex.
 While these methods can often achieve high image quality through a large number of iterations, they are time-consuming, and their convergence analyses were mostly not considered.
-In this article, we aim to develop an image reconstruction algorithm of the one-step type that takes into account both accuracy and efficiency, has mathematical guarantee, and is suitable for the geometric-inconsistency MSCT.
+In addition, many studies also consider the problem as solving a large-scale system of nonlinear equations [37, 16].
+In this article, we aim to develop an image reconstruction algorithm of the one-step type that takes into account both accuracy and efficiency, has mathematical guarantee, and is suitable for the geometric-inconsistent MSCT.
 The rest of this paper is organized as follows.
 In section 2, we introduce the forward data models of MSCT.
-In section 3, we propose a novel algorithm to solve the image reconstruction problem in geometric-inconsistency MSCT.
+In section 3, we propose a novel algorithm to solve the image reconstruction problem in geometric-inconsistent MSCT.
 We then prove the convergence of the proposed algorithm in section 4.
@@
 $\int_{0}^{E_{\text{max}}}s^{[q]}(E)\mathrm{d}E=1,$
-and $\mu(\boldsymbol{y},E)$ represents the energy-dependent LAC of interest at spatial position $\boldsymbol{y}\in\mathbb{R}^{n}$ ($n=2$ or $3$) and energy $E\in[0,E_{\text{max}}]$, and $\mathbb{R}^{n}$ the real Euclidean space, $E_{\text{max}}$ the maximum energy in the scan.
+and $\mu(\boldsymbol{y},E)$ represents the energy-dependent LAC of interest at spatial position $\boldsymbol{y}\in\mathbb{R}^{n}$ ($n=2$ or 333) and energy $E\in[0,E_{\text{max}}]$, and $\mathbb{R}^{n}$ the real Euclidean space, $E_{\text{max}}$ the maximum energy in the scan [7, 21, 2].
 The function $\mu$ of LAC is decomposed often into a linear combination of basis images $\{f_{d}\}_{d=1}^{D}$ as
 $\mu(\boldsymbol{y},E)=\sum_{d=1}^{D}b_{d}(E)f_{d}(\boldsymbol{y}),$ (2)
-where $b_{d}(E)$ denotes the expansion coefficient, and $D$ the number of basis images.
-For example, for the typical X-ray energy range in diagnostic DECT, $D=2$ is chosen often as photoelectric effect and Compton scatter are the dominant factors contributing to $\mu$; $b_{d}$ denotes the known function of mass attenuation coefficient (MAC) of the $d$-th basis material (e.g., water and bone); and $f_{d}$ represents the material-equivalent density function of the $d$-th basis material.
+where $b_{d}(E)$ denotes the expansion coefficient, and $D$ the number of basis images [4, 21].
+For example, for the typical X-ray energy range in diagnostic DECT, $D=2$ is chosen often as photoelectric effect and Compton scatter are the dominant factors contributing to $\mu$ in the physical effect decomposition model.
+For the basis material decomposition model, $b_{d}$ denotes the known function of mass attenuation coefficient (MAC) of the $d$-th basis material (e.g., water and bone); and $f_{d}$ represents the material-equivalent density function of the $d$-th basis material.
+Importantly, these two models are physically equivalent [21].
+Consequently, using two basis materials is sufficient to accurately describe the linear attenuation characteristics.
 Substituting 2 into 1, we obtain
 $g_{L(\theta^{[q]})}=\ln\int_{0}^{E_{\text{max}}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)\int_{L(\theta^{[q]})}f_{d}(\boldsymbol{y})\mathrm{d}l\right)\mathrm{d}E.$ (3)
-The formula in 3 is referred to as the CC-data model because $\theta^{[q]}$, $E$, and $\boldsymbol{y}$ are continuous variables.
+The formula in 3 is referred to as the CC-data model because $\theta^{[q]}$, $E$, and $\boldsymbol{y}$ are continuous variables [9].
 2.2 DD-data model
 In practice, data can be detected only for discrete energies and rays, and an image is reconstructed only on an array of discrete pixels/voxels.
-Hence, it is possible to develop a DD-data model, which may be constructed by implementing a specific discrete scheme into the CC-data model in 3.
+Hence, it is possible to develop a DD-data model, which may be constructed by implementing a specific discrete scheme into the CC-data model in 3 [9].
 However, in general, there is no direct connection between the two models.
-Assume that the energy domain $[0,E_{\text{max}}]$ is divided into $M$ intervals with equal size $\Delta_{E}$; the data $g_{j}^{[q]}$ are detected for discrete directions $\theta_{j}^{[q]}$ with $1\leq j\leq J_{q}$ under the $q$-th spectrum; each basis image $f_{d}$ is discretized by an array of size $I$.
+Assume that the energy domain $[0,E_{\text{max}}]$ is divided into $M$ intervals with equal size $\Delta_{E}$; the data $g_{j}^{[q]}$ are detected for discrete directions $\theta_{j}^{[q]}$ with $j=1,\dots,J_{q}$ under the $q$-th spectrum; each basis image $f_{d}$ is discretized by an array of size $I$.
 Then a DD-data model is obtained as
-$g_{j}^{[q]}=\ln\sum_{m=1}^{M}s^{[q]}_{m}\exp\left(-\sum_{d=1}^{D}b_{dm}\sum_{i=1}^{I}p^{[q]}_{ji}f_{di}\right)\quad\text{for $1\leq j\leq J_{q}$,}$ (4)
+$g_{j}^{[q]}=\ln\sum_{m=1}^{M}s^{[q]}_{m}\exp\left(-\sum_{d=1}^{D}b_{dm}\sum_{i=1}^{I}p^{[q]}_{ji}f_{di}\right)\quad\text{for $j=1,\dots,J_{q}$,}$ (4)
 where $s^{[q]}_{m}=s^{[q]}(m\Delta_{E})\Delta_{E}$ satisfying $\sum_{m=1}^{M}s^{[q]}_{m}=1$, $b_{dm}=b_{d}(m\Delta_{E})\Delta_{E}$, $f_{di}$ denotes the gray value of the $d$-th discrete basis image at voxel $i$, and $p^{[q]}_{ji}$ is often taken as the intersection length of ray $j$ with pixel/voxel $i$ of the image array [11].
@@
 Here we give the MSCT image reconstruction problem in continuous setting based on the CC-data model 3, and then present the corresponding algorithm.
+Before that, we introduce the required notation.
+Let $\mathcal{X}_{1},\ldots,\mathcal{X}_{D}$ be normed vector spaces with norms $\|\cdot\|_{\mathcal{X}_{1}},\ldots,\|\cdot\|_{\mathcal{X}_{D}}$, respectively.
+Their Cartesian product space is given by
+[…]
+The norm of an element $\boldsymbol{u}=[u_{1},\ldots,u_{D}]^{\boldsymbol{\mathsf{T}}}$ in this Cartesian product space can be defined as
+$\|\boldsymbol{u}\|_{\mathbfcal{X}^{D}}:=\left(\sum_{d=1}^{D}\|u_{d}\|^{2}_{\mathcal{X}_{d}}\right)^{1/2}.$
+Let $H^{\alpha}(\mathbb{R}^{n})$ be the Sobolev space of order $\alpha$ on $\mathbb{R}^{n}$, and $\Omega$ be a sufficiently regular open bounded subset in $\mathbb{R}^{n}$.
+Denote $H_{0}^{\alpha}(\Omega)=\{u\in H^{\alpha}(\mathbb{R}^{n})|\text{supp}(u)\subset\Omega\}$, where $\text{supp}(u):=\overline{\{\boldsymbol{y}\in\mathbb{R}^{n}|u(\boldsymbol{y})\neq 0\}}$ is the support of $u$.
 3.1.1 Continuous MSCT image reconstruction problem
-For simplicity, define $\mathbb{T}^{n}:=\{(\boldsymbol{\omega},\boldsymbol{x})|\boldsymbol{\omega}\in\mathbb{S}^{n-1},\boldsymbol{x}\in\boldsymbol{\omega}^{\bot}\}$, and
-$\mathcal{P}f(\boldsymbol{\omega},\boldsymbol{x}):=\int_{-\infty}^{+\infty}f(\boldsymbol{x}+t\boldsymbol{\omega})\mathrm{d}t,\quad(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n},$ (5)
+For simplicity, define $\mathbb{T}^{n}_{q}:=\{(\boldsymbol{\omega},\boldsymbol{x})|\boldsymbol{\omega}\in\mathbb{S}^{n-1}_{q}\subseteq\mathbb{S}^{n-1},\boldsymbol{x}\in\Omega_{q}\subseteq\boldsymbol{\omega}^{\bot}\}\subseteq\{(\boldsymbol{\omega},\boldsymbol{x})|\boldsymbol{\omega}\in\mathbb{S}^{n-1},\boldsymbol{x}\in\boldsymbol{\omega}^{\bot}\}=:\mathbb{T}^{n}$ as the set of scanning geometric parameters under the $q$-th energy spectrum, where $q=1,\ldots,Q$.
+Here $\mathbb{S}^{n-1}$ represents the unit sphere in $\mathbb{R}^{n}$ and $\boldsymbol{\omega}^{\bot}$ denotes the hyperplane perpendicular to $\boldsymbol{\omega}$.
+Let
+$\mathcal{P}_{q}f(\boldsymbol{\omega},\boldsymbol{x}):=\int_{-\infty}^{+\infty}f(\boldsymbol{x}+t\boldsymbol{\omega})~\mathrm{d}t,\quad(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q},$ (5)
 which is the X-ray transform of a certain function $f$ defined on $\mathbb{R}^{n}$, equivalently the Radon transform if $n=2$ [31].
-Here $\mathbb{S}^{n-1}$ represents the unit sphere in $\mathbb{R}^{n}$, and $\boldsymbol{\omega}^{\bot}$ denotes the hyperplane perpendicular to $\boldsymbol{\omega}$.
-As the X-ray $l(\theta^{[q]})$ in 3 can be parametrized by $(\boldsymbol{\omega},\boldsymbol{x})$, we rewrite 5 as
-$\mathcal{P}f(\boldsymbol{\omega},\boldsymbol{x})=\int_{l(\theta^{[q]})}f(\boldsymbol{y})\mathrm{d}l.$ (6)
+Note that if $\mathbb{T}^{n}_{1}=\cdots=\mathbb{T}^{n}_{Q}$, then $\mathcal{P}_{1}=\cdots=\mathcal{P}_{Q}$, this situation is called geometric consistency, otherwise it is called geometric inconsistency.
+For each energy spectrum $q$, denote the X-ray transform in 5 by the mapping
+$\mathcal{P}_{q}:\mathcal{F}=H_{0}^{-1/2}(\Omega)\longrightarrow\tilde{\mathcal{G}}_{q}=\mathcal{P}_{q}(\mathcal{F})\subseteq\mathcal{L}^{2}(\mathbb{T}^{n}_{q})=\mathcal{G}_{q}.$
+As the X-ray $L(\theta^{[q]})$ in 3 can be parametrized by $(\boldsymbol{\omega},\boldsymbol{x})$, we translate 5 into
+$\mathcal{P}_{q}f(\boldsymbol{\omega},\boldsymbol{x})=\int_{L(\theta^{[q]})}f(\boldsymbol{y})~\mathrm{d}l.$ (6)
 Using 6, the right-hand side of 3 can be rewritten as
-$K^{[q]}(\boldsymbol{f}):=\ln\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)\mathrm{d}E,$ (7)
+$K^{[q]}(\boldsymbol{f}):=\ln\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}f_{d}\right)\mathrm{d}E,$ (7)
 where $\boldsymbol{f}=[f_{1},\ldots,f_{D}]^{\boldsymbol{\mathsf{T}}}$.
-Consequently, the aim of the continuous MSCT image reconstruction is to reconstruct $\boldsymbol{f}$ from the measured data functions $\{g^{[q]}\}_{q=1}^{Q}$ such that
+Consequently, the aim of the continuous MSCT image reconstruction is to reconstruct $\boldsymbol{f}$ from the measured data $\boldsymbol{g}=[g^{[1]},\ldots,g^{[Q]}]^{\boldsymbol{\mathsf{T}}}$ such that
 $K^{[q]}(\boldsymbol{f})=g^{[q]}\quad\text{for}~q=1,\ldots,Q.$ (8)
@@
 $\boldsymbol{K}(\boldsymbol{f})=\boldsymbol{g},$ (9)
-where $\boldsymbol{K}(\boldsymbol{f})=[K^{[1]}(\boldsymbol{f}),\ldots,K^{[Q]}(\boldsymbol{f})]^{\boldsymbol{\mathsf{T}}}$ and $\boldsymbol{g}=[g^{[1]},\ldots,g^{[Q]}]^{\boldsymbol{\mathsf{T}}}$.
+where $\boldsymbol{K}(\boldsymbol{f})=[K^{[1]}(\boldsymbol{f}),\ldots,K^{[Q]}(\boldsymbol{f})]^{\boldsymbol{\mathsf{T}}}$.
 3.1.2 The proposed algorithm for the continuous problem
-Here we propose the algorithm AFIRE for solving the continuous MSCT image reconstruction problem.
-Before that, we introduce the required notation. Let $\mathcal{X}$ be a normed vector space with norm $\|\cdot\|_{\mathcal{X}}$, and
-$\mathcal{X}^{D}:=\underbrace{\mathcal{X}\times\cdots\times\mathcal{X}}_{D}$
-be the Cartesian product of the space $\mathcal{X}$ with the norm defined by
-$\|\boldsymbol{u}\|_{{\mathcal{X}}^{D}}:=\left(\sum_{d=1}^{D}\|u_{d}\|^{2}_{\mathcal{X}}\right)^{1/2},$
 […]
-Denote the X-ray transform in 5 by the mapping
-$\mathcal{P}:\mathcal{F}=H_{0}^{-1/2}(\Omega)\longrightarrow\tilde{\mathcal{G}}=\mathcal{P}(\mathcal{F})\subset\mathcal{L}^{2}(\mathbb{T}^{n})=\mathcal{G}.$
-By [30, theorem 2.18], it is known that $\mathcal{P}$ is invertible.
 First of all, we have the following proposition.
 Proposition 1.
-Let $\{s^{[q]}\}_{q=1}^{Q}$ be nonnegative normalized functions and $\{b_{d}\}_{d=1}^{D}$ be nonnegative bounded functions.
-Then for each $q$, the functional in 7 can be represented as a mapping $K^{[q]}:\mathcal{F}^{D}\rightarrow\mathcal{G}$ and is Gâteaux differentiable, and its Gâteaux derivative at $\boldsymbol{f}\in\mathcal{F}^{D}$ for $\boldsymbol{h}=[h_{1},\ldots,h_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathcal{F}^{D}$ is computed as
-$\mathcal{D}K^{[q]}(\boldsymbol{f},\boldsymbol{h})=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}h_{d}\right)\mathrm{d}E,$ (10)
+[…]
+$\mathcal{D}K^{[q]}(\boldsymbol{f},\boldsymbol{h})=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}h_{d}\right)\mathrm{d}E,$ (10)
 where
-$\Phi^{[q]}(\boldsymbol{f};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)\mathrm{d}E}.$ (11)
-Moreover, both the $K^{[q]}$ and $\boldsymbol{K}:\mathcal{F}^{D}\rightarrow\mathcal{G}^{Q}$ are Fréchet differentiable.
+$\Phi^{[q]}(\boldsymbol{f};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}f_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}f_{d}\right)\mathrm{d}E}.$ (11)
+Moreover, both the $K^{[q]}$ and $\boldsymbol{K}:\mathbfcal{F}^{D}\rightarrow\mathbfcal{G}^{Q}$ are Fréchet differentiable.
 Proof.
@@
 We start with analyzing the Gâteaux derivative of $K^{[q]}$ in 10.
-Notice that the $\Phi^{[q]}(\boldsymbol{f};E)$ in 11 is actually related to the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$ since the term $\mathcal{P}f_{d}$ depends on that parameter as in 6.
-For a certain real constant set $\{C_{d}\}_{d=1}^{D}$, if there exists one $\hat{\boldsymbol{f}}=[\hat{f}_{1},\ldots,\hat{f}_{D}]^{\boldsymbol{\mathsf{T}}}$ such that
-$\mathcal{P}\hat{f}_{d}(\boldsymbol{\omega},\boldsymbol{x})\equiv C_{d},\quad\forall(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n},$ (12)
-then $\Phi^{[q]}(\hat{\boldsymbol{f}};E)$ becomes to be independent of the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})$.
-Together with the linearity of $\mathcal{P}$, the Gâteaux derivative at $\hat{\boldsymbol{f}}$ for $\boldsymbol{h}$ can be rewritten as
-$\mathcal{D}K^{[q]}(\hat{\boldsymbol{f}},\boldsymbol{h})=-\mathcal{P}\left(\sum_{d=1}^{D}\varphi^{[q]}_{d}h_{d}\right).$ (13)
+Notice that the $\Phi^{[q]}(\boldsymbol{f};E)$ in 11 is actually related to the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q}$ since the term $\mathcal{P}_{q}f_{d}$ depends on that parameter as in 6.
+For a certain real constant set $C_{d}^{[q]}\}_{d=1}^{D}$, if there exists one $\hat{\boldsymbol{f}}=[\hat{f}_{1},\ldots,\hat{f}_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathbfcal{F}^{D}$ such that
+$\mathcal{P}_{q}\hat{f}_{d}(\boldsymbol{\omega},\boldsymbol{x})\equiv C_{d}^{[q]},\quad\forall(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q},$ (12)
+then $\Phi^{[q]}(\hat{\boldsymbol{f}};E)$ is independent of the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})$.
+Together with the linearity of $\mathcal{P}_{q}$, the Gâteaux derivative at $\hat{\boldsymbol{f}}$ for $\boldsymbol{h}$ can be rewritten as
+$\mathcal{D}K^{[q]}(\hat{\boldsymbol{f}},\boldsymbol{h})=-\mathcal{P}_{q}\left(\sum_{d=1}^{D}\varphi^{[q]}_{d}h_{d}\right).$ (13)
 Here
-$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E,$ (14)
+$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)~\mathrm{d}E,$ (14)
 where
-$\Phi^{[q]}(\hat{\boldsymbol{f}};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}\right)\mathrm{d}E}.$
-Obviously, perhaps the most simple example is taking $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathcal{F}^{D}$ for $C_{d}=0$, $d=1,\ldots,D$ that satisfies the condition of 12.
-There may be other points that satisfy this condition, but searching for them is beyond the scope of the current research.
+$\Phi^{[q]}(\hat{\boldsymbol{f}};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}^{[q]}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}^{[q]}\right)\mathrm{d}E}.$
+Remark 1.
+The simplest example is taking $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathbfcal{F}^{D}$ for all of the $C_{d}^{[q]}=0$, $d=1,\ldots,D$ and $q=1,\ldots,Q$ that satisfies the condition of 12.
+Although there may be other points that satisfy this condition, their explicit identification is not critical and is beyond the scope of the current research.
+This is because both the practical implementation and the theoretical convergence of the proposed algorithm are independent of the explicit form of such a point.
 Following the insight above, the Gâteaux derivative of $\boldsymbol{K}$ at the $\hat{\boldsymbol{f}}$ determined by 12 can be given as the mapping below
-$\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot):\mathcal{F}^{D}$ $\longrightarrow\tilde{\mathcal{G}}^{Q}$
+$\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot):\mathbfcal{F}^{D}$ $\longrightarrow\tilde{\mathbfcal{G}}^{Q}$
 $\boldsymbol{h}$ $\longmapsto-\boldsymbol{\mathcal{P}}(\boldsymbol{\varphi}\boldsymbol{h}),$ (15)
-[…]
+where $\boldsymbol{\mathcal{P}}=\operatorname*{\texttt{diag}}(\mathcal{P}_{1},\ldots,\mathcal{P}_{Q})$ denotes a diagonal operator composed of $Q$ X-ray transforms, and $\boldsymbol{\varphi}=\bigl{(}\varphi^{[q]}_{d}\bigr{)}_{Q\times D}$ represents the constant matrix with entries $\varphi^{[q]}_{d}$ given in 14.
 To proceed, we study the invertibility of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$. We have the following result.
 Proposition 2.
-Assume that the conditions in proposition 1 hold, and the matrix $\boldsymbol{\varphi}$ is invertible. Then the $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ in section 3.1.2 is invertible.
+Assume that $Q=D$, the conditions in proposition 1 hold, and the X-ray transform $\mathcal{P}_{q}:\mathcal{F}\rightarrow\tilde{\mathcal{G}}_{q}$ for each $q$ and the matrix $\boldsymbol{\varphi}$ are invertible.
+Then the $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ in section 3.1.2 is invertible.
 Proof.
-Since $\mathcal{P}$ is invertible from $\mathcal{F}$ to $\tilde{\mathcal{G}}$, $\boldsymbol{\mathcal{P}}$ is invertible from $\mathcal{F}^{D}$ to $\tilde{\mathcal{G}}^{Q}$ when $Q=D$.
+Since $\mathcal{P}_{1},\ldots,\mathcal{P}_{Q}$ are invertible, $\boldsymbol{\mathcal{P}}$ is invertible.
 Combining with the invertibility of $\boldsymbol{\varphi}$, we obtain the invertibility of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$.
@@
 Remark 2.
-The invertibility of $\boldsymbol{\varphi}$ implies $Q=D$, which means that the number of used energy spectra should be equal to the number of basis images.
-Actually, this necessary condition is commonly satisfied in clinical applications, for example, $Q=D=2$ in DECT.
-Moreover, denote $\mathcal{P}^{-1}$ as the inverse of $\mathcal{P}$, and naturally
-$\boldsymbol{\mathcal{P}}^{-1}=\operatorname*{\texttt{diag}}(\mathcal{P}^{-1},\ldots,\mathcal{P}^{-1})$
+If $\mathbb{S}^{n-1}_{q}\subseteq\mathbb{S}^{n-1}$ meets every equatorial circle of $\mathbb{S}^{n-1}$, it follows from [32, theorem 2.11] that $\mathcal{P}_{q}:\mathcal{F}\rightarrow\tilde{\mathcal{G}}_{q}$ is invertible.
+The condition of $Q=D$ means that the number of used energy spectra should be equal to the number of basis images, which is commonly satisfied in clinical applications.
+For example, $Q=D=2$ in DECT.
+Moreover, denote $\mathcal{P}^{-1}_{q}$ as the inverse of $\mathcal{P}_{q}$, and naturally
+$\boldsymbol{\mathcal{P}}^{-1}=\operatorname*{\texttt{diag}}(\mathcal{P}_{1}^{-1},\ldots,\mathcal{P}_{Q}^{-1})$
 is the inverse of $\boldsymbol{\mathcal{P}}$. From proposition 2, we can obtain the inverse mapping of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ as
-$(\mathcal{D}\boldsymbol{K})^{-1}(\hat{\boldsymbol{f}},\cdot):\tilde{\mathcal{G}}^{Q}$ $\longrightarrow\mathcal{F}^{D}$
+$(\mathcal{D}\boldsymbol{K})^{-1}(\hat{\boldsymbol{f}},\cdot):\tilde{\mathbfcal{G}}^{Q}$ $\longrightarrow\mathbfcal{F}^{D}$
 $\boldsymbol{r}$ $\longmapsto-\boldsymbol{\varphi}^{-1}\boldsymbol{\mathcal{P}}^{-1}(\boldsymbol{r}).$ (16)
-Hence, by leveraging the simplified Newton method [10], the proposed iterative scheme for solving 9 can be formulated as
+Hence, by leveraging the simplified Newton method [12, 33], the proposed iterative scheme for solving 9 can be formulated as
 $\boldsymbol{f}^{k+1}=\boldsymbol{f}^{k}-\boldsymbol{\varphi}^{-1}\boldsymbol{\mathcal{P}}^{-1}\bigl{(}\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k})\bigr{)}.$ (17)
-Note that 17 is well-defined if $\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k})\in\tilde{\mathcal{G}}^{Q}$ for $\boldsymbol{f}^{k}\in\mathcal{F}^{D}$.
+Note that 17 is well-defined if $\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k})\in\tilde{\mathbfcal{G}}^{Q}$ for $\boldsymbol{f}^{k}\in\mathbfcal{F}^{D}$, and the standard simplified Newton method considers the derivative at the initial point $\boldsymbol{f}^{0}$.
 Based on the iterative scheme in 17, we present the algorithm AFIRE for solving the continuous MSCT image reconstruction problem in algorithm 1.
 Algorithm 1 The proposed algorithm for solving the continuous problem in 8 (or 9).
-1:Initialize: Given the measured data $\{g^{[q]}\}_{q=1}^{Q}$, the energy spectra distributions $\{s^{[q]}\}_{q=1}^{Q}$, the MAC functions $\{b_{d}\}_{d=1}^{D}$ and a certain real constant set $\{C_{d}\}_{d=1}^{D}$.
+1:Initialize: Given $Q=D$, the measured data $\{g^{[q]}\}_{q=1}^{Q}$, the distributions of energy spectra $\{s^{[q]}\}_{q=1}^{Q}$, the MAC functions $\{b_{d}\}_{d=1}^{D}$ and a certain real constant set $\{C_{d}\}_{d=1}^{D}$.
 Initialize the basis image $\boldsymbol{f}^{0}$.
 Let $k\leftarrow 0$.
-[…]
+2:Compute the constant matrix $\boldsymbol{\varphi}=(\varphi^{[q]}_{d})_{Q\times D}$ by 14 and the inverse $\boldsymbol{\varphi}^{-1}$.
 3:Loop:
 4: Compute the residual
-$\begin{bmatrix}\varLambda_{1}^{k}\\ \vdots\\ \varLambda_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}g^{[1]}-K^{[1]}(\boldsymbol{f}^{k})\\ \vdots\\ g^{[Q]}-K^{[Q]}(\boldsymbol{f}^{k})\end{bmatrix}.$
+$\begin{bmatrix}\varLambda_{1}^{k}\\ \vdots\\ \varLambda_{Q}^{k}\end{bmatrix}\longleftarrow\begin{bmatrix}g^{[1]}-K^{[1]}(\boldsymbol{f}^{k})\\ \vdots\\ g^{[Q]}-K^{[Q]}(\boldsymbol{f}^{k})\end{bmatrix}.$
 5: Update $\boldsymbol{f}^{k+1}$ by
-$\begin{bmatrix}f_{1}^{k+1}\\ \vdots\\ f_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}f_{1}^{k}\\ \vdots\\ f_{D}^{k}\end{bmatrix}-\boldsymbol{\varphi}^{-1}\begin{bmatrix}\mathcal{P}^{-1}\varLambda_{1}^{k}\\ \vdots\\ \mathcal{P}^{-1}\varLambda_{Q}^{k}\end{bmatrix}.$
+$\begin{bmatrix}f_{1}^{k+1}\\ \vdots\\ f_{D}^{k+1}\end{bmatrix}\longleftarrow\begin{bmatrix}f_{1}^{k}\\ \vdots\\ f_{D}^{k}\end{bmatrix}-\boldsymbol{\varphi}^{-1}\begin{bmatrix}\mathcal{P}^{-1}_{1}\varLambda_{1}^{k}\\ \vdots\\ \mathcal{P}^{-1}_{Q}\varLambda_{Q}^{k}\end{bmatrix}.$
 6: If some given termination condition is satisfied, then output $\boldsymbol{f}^{k+1}$;
@@
 Remark 3.
-In algorithm 1, one can specifically choose the $C_{d}=0$ for $d=1,\ldots,D$ with $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathcal{F}^{D}$ to fulfill the condition in 12.
+In algorithm 1, one can specifically choose the $C_{d}^{[q]}=0$ for $d=1,\ldots,D$ and $q=1,\ldots,Q$ with $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathbfcal{F}^{D}$ to fulfill the condition of 12.
 The corresponding formula of 14 becomes
-$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}s^{[q]}(E)b_{d}(E)\mathrm{d}E.$
+$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}s^{[q]}(E)b_{d}(E)~\mathrm{d}E.$
 3.2 The discrete setting
-Here, we start from the DD-data model 4 and then present the associated algorithm for solving the geometric-inconsistency full-scan MSCT image reconstruction problem.
+Here, we start from the DD-data model 4 and then present the associated algorithm for solving the geometric-inconsistent MSCT image reconstruction problem under the mildly full-scan case.
 3.2.1 Discrete MSCT image reconstruction problem
-As described in section 2.2, for the discrete case, the forward operator of ray $j$ with the $q$-th energy spectrum may be formulated as
+As described in section 2.2, for the discrete case, the forward operator of ray $j$ under the $q$-th energy spectrum may be formulated as
 $K_{j}^{[q]}(\mathbf{f})=\ln\sum_{m=1}^{M}s^{[q]}_{m}\exp\left(-\sum_{d=1}^{D}b_{dm}\sum_{i=1}^{I}p^{[q]}_{ji}f_{di}\right).$ (18)
@@
 Hence, the discrete MSCT image reconstruction problem is to determine $\mathbf{f}$ from the measured data $\{g_{j}^{[q]}\}_{j,q=1}^{J_{q},Q}$ by the system of nonlinear equations as below
-$K_{j}^{[q]}(\mathbf{f})=g_{j}^{[q]}\quad\text{for}~{}q=1,\ldots,Q~{}\text{and}~{}j=1,\ldots,J_{q}.$ (19)
-The compact form of 19 can be expressed as
+$K_{j}^{[q]}(\mathbf{f})=g_{j}^{[q]}\quad\text{for}~j=1,\ldots,J_{q}~\text{and}~q=1,\ldots,Q.$ (19)
+Here $J_{q}$ is the number of rays under the $q$-th energy spectrum. The compact form of 19 can be expressed as
 $\mathbf{K}(\mathbf{f})=\mathbf{g}.$ (20)
@@
 $\mathbf{K}(\mathbf{f})=\begin{bmatrix}\mathbf{K}^{[1]}(\mathbf{f})\\ \vdots\\ \mathbf{K}^{[Q]}(\mathbf{f})\end{bmatrix},\mathbf{K}^{[q]}(\mathbf{f})=\begin{bmatrix}K_{1}^{[q]}(\mathbf{f})\\ \vdots\\ K_{J_{q}}^{[q]}(\mathbf{f})\end{bmatrix},\mathbf{g}=\begin{bmatrix}\mathbf{g}^{[1]}\\ \vdots\\ \mathbf{g}^{[Q]}\end{bmatrix},\mathbf{g}^{[q]}=\begin{bmatrix}g^{[q]}_{1}\\ \vdots\\ g^{[q]}_{J_{q}}\end{bmatrix}.$
+[…]
+Remark 4.
+In the geometric-consistent case, the same scanning geometries across different energy spectra allow 20 to be divided into a lot of very small-scale independent subsystems along each X-ray path, enabling high solving efficiency via the two-step DDD method [18].
+In contrast, the geometric inconsistency prevents this kind of division, requiring to directly solve a large-scale nonlinear system with $\sum_{q=1}^{Q}J_{q}$ equations (e.g., 294,912 equations considered in section 5.2).
+If the special structure is ignored, solving such a large-scale nonlinear problem with high precision is quite time-consuming.
 3.2.2 The proposed algorithm for the discrete problem
@@
 $\mathbf{w}_{j}^{[q]}(\mathbf{f})=\frac{\mathbf{s}^{[q]}\odot\exp\bigl{(}\mathbf{z}_{j}^{[q]}(\mathbf{f})\bigr{)}}{(\mathbf{s}^{[q]})^{\boldsymbol{\mathsf{T}}}\exp\bigl{(}\mathbf{z}_{j}^{[q]}(\mathbf{f})\bigr{)}},$ (22)
-and $\odot$ denotes the operation of the element-wise product.
+and $\odot$ denotes the operation of the elementwise product.
 Our insight
@@
 $\hat{\mathbf{w}}^{[q]}:=\mathbf{w}_{j}^{[q]}(\hat{\mathbf{f}}).$
-Note that probably the most simple example satisfying the condition of 23 is taking $\hat{\mathbf{f}}=\boldsymbol{0}$, which leads to $C_{d}^{[q]}=0$ for any indices $q$ and $d$.
-But searching for the other points is beyond the scope of this work.
+Remark 5.
+The simplest example that satisfying the condition of 23 is taking $\hat{\mathbf{f}}=\boldsymbol{0}$, which leads to $C_{d}^{[q]}=0$ for any $q$ and $d$.
+Similarly to remark 1, the explicit identification of other such points is not critical and is beyond the scope of this work.
 Considering 21 at the $\hat{\mathbf{f}}$ in 23, we obtain that
@@
 where $\mathbf{P}_{q}=[\mathbf{p}_{1}^{[q]},\dots,\mathbf{p}_{J_{q}}^{[q]}]^{\boldsymbol{\mathsf{T}}}$ is the projection matrix under the $q$-th energy spectrum.
+Note that $\mathbf{P}_{1}=\cdots=\mathbf{P}_{Q}$ in the geometric-consistent case, but it fails to hold in the geometric-inconsistent case.
 Furthermore, the associated Jacobian matrix of $\mathbf{K}$ at $\hat{\mathbf{f}}$ is
@@
 where $\mathbf{P}^{-1}=\operatorname*{\texttt{diag}}(\mathbf{P}_{1}^{-1},\ldots,\mathbf{P}_{Q}^{-1})$. ∎
-Again, by leveraging the simplified Newton method as before, the proposed iterative scheme for solving the discrete problem in 19 (or 20) can be formulated as
+Again, by leveraging the simplified Newton method [12, 33], the proposed iterative scheme for solving the discrete problem in 19 (or 20) can be formulated as
 $\mathbf{f}^{k+1}=\mathbf{f}^{k}-\boldsymbol{\phi}^{-1}\otimes\mathbf{P}^{-1}\bigl{(}\mathbf{g}-\mathbf{K}(\mathbf{f}^{k})\bigr{)}.$ (27)
+Note that the standard simplified Newton method utilizes the Jacobian matrix at the initial point $\mathbf{f}^{0}$.
 More clearly, based on 27, we present the proposed algorithm AFIRE for solving 19 (or 20) in algorithm 2, which may also be seen as a specific discretized counterpart of algorithm 1.
 Algorithm 2 The proposed algorithm for solving the discrete problem in 19 (or 20).
-1:Initialize: Given the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$.
+1:Initialize: Given $Q=D$, the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$.
 Initialized the basis images $\{\mathbf{f}_{d}^{0}\}_{d=1}^{D}$.
@@
 4: Compute the residual
-$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
+$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\longleftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
 5: Update $\mathbf{f}^{k+1}$ by
-$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{1}^{-1}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{Q}^{-1}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$
+$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\longleftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{1}^{-1}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{Q}^{-1}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$ (28)
 6: If some given termination condition is satisfied, then output $\mathbf{f}^{k+1}$;
@@
 $\phi^{[q]}_{d}=\sum_{m=1}^{M}b_{dm}s^{[q]}_{m}.$
-As a matter of fact, the inverse of the matrix $\mathbf{P}$ in 27 is difficult to calculate accurately, as all of its diagonal elements $\mathbf{P}_{1},\ldots,\mathbf{P}_{Q}$ may not even be invertible square matrices, even if they are, the computational cost of calculating the inverse of a large-scale matrix is very high.
-Therefore, we can only obtain an approximation for the inverse of $\mathbf{P}$, and the actual iterative scheme used in algorithm 2 should be
+Remark 7.
+As a matter of fact, the inverse of the matrix $\mathbf{P}$ in 27 is difficult to calculate accurately.
+That is because all of its diagonal elements $\mathbf{P}_{1},\ldots,\mathbf{P}_{Q}$ may not even be invertible square matrices, even if they are, the computational cost of calculating the inverse of a large-scale matrix is very high.
+Therefore, we can only obtain an approximation for the inverse of $\mathbf{P}$, and the actual iterative scheme used for 27 should be
 $\mathbf{f}^{k+1}=\mathbf{f}^{k}-\boldsymbol{\phi}^{-1}\otimes\mathbf{P}^{+}\bigl{(}\mathbf{g}-\mathbf{K}(\mathbf{f}^{k})\bigr{)},$ (29)
@@
 Here $\mathbf{P}_{q}^{+}$ denotes the approximation for the inverse of $\mathbf{P}_{q}$.
-Subsequently, we present the actual algorithm AFIRE used for solving the discrete problem of 19 (or 20) in algorithm 3.
-Algorithm 3 The actual algorithm used for solving the discrete problem 19 (or 20).
-1:Initialize: Given the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$.
-Initialized the basis images $\{\mathbf{f}_{d}^{0}\}_{d=1}^{D}$.
-Let $k\leftarrow 0$.
-2:Compute the constant matrix $\boldsymbol{\phi}=\bigl{(}\phi^{[q]}_{d}\bigr{)}_{Q\times D}$ from 25 and the inverse $\boldsymbol{\phi}^{-1}$.
-3:Loop:
-4: Compute the residual
-$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
-5: Update $\mathbf{f}^{k+1}$ by
-$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{\!1}^{+}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{\!Q}^{+}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$
-6: If some given termination condition is satisfied, then output $\mathbf{f}^{k+1}$;
-otherwise, let $k\leftarrow k+1$, goto Loop.
+The update of 28 should be replaced with
+$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\longleftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{1}^{+}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{Q}^{+}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$ (30)
 Remark 8.
@@
 4 Convergence analysis
-In this section, we study the convergence of the proposed algorithms that given in algorithms 1, 2 and 3.
+In this section, we study the convergence of the proposed algorithms that given in algorithms 1 and 2.
 Assuming that both the solutions exist to the continuous problem 9 and the discrete one 20, and we refer to $\boldsymbol{f}^{*}$ and $\mathbf{f}^{*}$ as the solutions, respectively.
@@
 If $\sigma_{\min}(\cdot)$ is nonvanishing, define the scaled condition number by $\kappa_{F}(\cdot):=\|\cdot\|_{F}/\sigma_{\min}(\cdot)$.
-Denote $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ as the $1$-norm and Euclidean norm of a vector or the spectral norm of a matrix, respectively.
+Denote $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ as the 111-norm and Euclidean norm of a vector or the spectral norm of a matrix, respectively.
 4.1 Convergence analysis for the continuous setting
@@
 Lemma 1.
-Suppose that the conditions in proposition 1 hold and $\boldsymbol{\varphi}$ is invertible.
+Suppose that the conditions in proposition 2 hold, and $\mathbb{T}^{n}_{q}=\mathbb{T}^{n}$ for $q=1,\ldots,Q$.
 Assume that there exists a constant $\eta>0$ such that, for any nonnegative normalized function $\rho$ and any indices $q,d$,
 $\Bigl{|}\int_{0}^{E_{\max}}\left(\rho(E)-\Phi^{[q]}(\hat{\boldsymbol{f}};E)\right)b_{d}(E)\mathrm{d}E\Bigr{|}\leq\eta\Bigl{|}\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E\Bigr{|}.$
-Then, for any $\boldsymbol{f},\boldsymbol{h}\in\mathcal{F}^{D}$,
-$\|\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}.$ (29)
+Then, for any $\boldsymbol{f},\boldsymbol{h}\in\mathbfcal{F}^{D}$,
+$\|\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathbfcal{G}^{Q}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathbfcal{G}^{Q}}.$ (31)
 Proof.
@@
 Suppose that the conditions in lemma 1 hold and $\eta\kappa_{F}(\boldsymbol{\varphi})<1$, and the iterative scheme 17 is well-defined.
-Then, for any initial point $\boldsymbol{f}^{0}\in\mathcal{F}^{D}$, the iteration sequence $\{\boldsymbol{f}^{k}\}$ generated by algorithm 1 converges to the unique solution of 9.
+Then, for any initial point $\boldsymbol{f}^{0}\in\mathbfcal{F}^{D}$, the iteration sequence $\{\boldsymbol{f}^{k}\}$ generated by algorithm 1 converges to the unique solution of 9.
 Proof.
@@
 The following estimate is obtained
-$\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},$ $\boldsymbol{f}^{k+1}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}$
-$=\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k+1}-\boldsymbol{f}^{k})+\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}$
+$\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},$ $\boldsymbol{f}^{k+1}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}$
+$=\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k+1}-\boldsymbol{f}^{k})+\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}$
 […]
-$=\Big{\|}\int_{0}^{1}\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\mathrm{d}t\Big{\|}_{\mathcal{G}^{D}}$
-$\leq\int_{0}^{1}\|\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}\mathrm{d}t$
+$=\Big{\|}\int_{0}^{1}\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\mathrm{d}t\Big{\|}_{\mathbfcal{G}^{Q}}$
+$\leq\int_{0}^{1}\|\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}\mathrm{d}t$
 […]
 With $\eta\kappa_{F}(\boldsymbol{\varphi})<1$, we have
-$\big{\|}\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\big{\|}_{\mathcal{G}^{D}}\longrightarrow 0\quad\text{as}~{}k\longrightarrow\infty.$
+$\big{\|}\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\big{\|}_{\mathbfcal{G}^{Q}}\longrightarrow 0\quad\text{as}~k\longrightarrow\infty.$
 Based on proposition 2, we know that $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ is invertible. Then, we obtain
-$\|\boldsymbol{f}^{k}-\boldsymbol{f}^{*}\|_{\mathcal{F}^{D}}\longrightarrow 0\quad\text{as}~{}k\longrightarrow\infty.$
+$\|\boldsymbol{f}^{k}-\boldsymbol{f}^{*}\|_{\mathbfcal{F}^{D}}\longrightarrow 0\quad\text{as}~k\longrightarrow\infty.$
 This completes the proof of the convergence.
 Finally, we prove the uniqueness of the solution. If there exists another solution $\bar{\boldsymbol{f}}^{*}$ satisfies 9, according to the above estimate, we have
-$\|\boldsymbol{K}(\bar{\boldsymbol{f}}^{*})-\boldsymbol{K}(\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}.$
-Thus, $\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}=0$, which implies that $\bar{\boldsymbol{f}}^{*}=\boldsymbol{f}^{*}$. ∎
+$\|\boldsymbol{K}(\bar{\boldsymbol{f}}^{*})-\boldsymbol{K}(\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}.$
+Thus, $\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}=0$, which implies that $\bar{\boldsymbol{f}}^{*}=\boldsymbol{f}^{*}$. ∎
 4.2 Convergence analysis for the discrete setting
@@
 See section A.3 for the proof. ∎
-Consider the scheme 28 and 26, and denote the residual in each iteration by
+Considering the iterative scheme 29, and using 26, we denote a residual at each iteration by
 $\boldsymbol{\delta}^{k}=\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{k})-\bigl{(}\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\bigr{)}.$ (33)
@@
 $\gamma=\tilde{\eta}\kappa_{F}(\mathbf{P}\otimes\boldsymbol{\phi}).$
-Under proper conditions, we have the following global convergence of algorithm 3.
+Under proper conditions, we have the following global convergence of the iterative scheme 29.
 Theorem 2.
-Assume that the conditions in lemma 2 hold and $\gamma<1$. The solution to 20 is unique.
-Moreover, for any $k$, $\|\boldsymbol{\delta}^{k}\|_{2}\leq\zeta\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2}$, where $\zeta<(1-\gamma)/(1+\gamma)$.
-Then, the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 28 converges to the unique solution of 20.
+Assume that the conditions in lemma 2 hold and $\gamma<1$. Then, the solution to 20 is unique.
+Moreover, suppose that for any $k$, the residual in 33 satisfies
+$\|\boldsymbol{\delta}^{k}\|_{2}\leq\zeta\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2},$ (34)
+where the constant $\zeta<(1-\gamma)/(1+\gamma)$. Then, the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 29 converges to the unique solution of 20.
 Proof.
-By mean value theorem and lemma 2, for any $\mathbf{f}$, we have the following estimate
+By the mean value theorem and lemma 2, for any $\mathbf{f}$, we have the following estimate
 $\|\mathbf{K}(\mathbf{f})-\mathbf{K}(\mathbf{f}^{*})-$ $\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\|_{2}$
@@
 $\leq\gamma\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\|_{2}.$ (35)
-Since $\gamma<1$ and $\nabla\mathbf{K}(\hat{\mathbf{f}})$ has full column rank, let the $\mathbf{f}$ in section 4.2 be another solution if it exists, the uniqueness of the solution is immediately obtained by section 4.2.
+Letting $\bar{\mathbf{f}}^{*}$ be another solution if it exists, by section 4.2, we obtain
+$\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\bar{\mathbf{f}}^{*}-\mathbf{f}^{*})\|_{2}\leq\gamma\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\bar{\mathbf{f}}^{*}-\mathbf{f}^{*})\|_{2}.$
+Using the assumption of $\gamma<1$, it follows that
+$\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\bar{\mathbf{f}}^{*}-\mathbf{f}^{*})\|_{2}=0.$
+Since $\nabla\mathbf{K}(\hat{\mathbf{f}})$ has full column rank, we conclude that $\bar{\mathbf{f}}^{*}=\mathbf{f}^{*}$. Thus, the solution to 20 is unique.
 Furthermore, from section 4.2, the triangle inequality implies that
 $\|\mathbf{K}(\mathbf{f})-\mathbf{K}(\mathbf{f}^{*})\|_{2}\leq(1+\gamma)\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}-\mathbf{f}^{*})\|_{2}.$
-Hence, by 31 and 4.2, and the above inequality in order, we have
+Hence, by 33, 34 and 4.2, and the above inequality in order, we have
 $\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{*})\|_{2}$ $=\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{k})+\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}$
@@
 $\leq\gamma\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}+\zeta\|\mathbf{K}(\mathbf{f}^{k})-\mathbf{g}\|_{2}$
-$\leq\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}.$
+$\leq\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}.$ (36)
 It follows from $\gamma+\zeta(1+\gamma)<1$ that the sequence $\{\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}\}$ converges to zero.
@@
 ∎
-For the scheme 27, the residual $\boldsymbol{\delta}^{k}$ defined in 31 is vanishing. Following theorem 2, it is easy to obtain the convergence of algorithm 2.
+Note that for the iterative scheme 27, the residual $\boldsymbol{\delta}^{k}$ in 33 is always vanishing.
+Following theorem 2, it is easy to obtain the convergence of the scheme 27, namely, algorithm 2.
+We present this result by the following corollary.
 Corollary 1.
 Assume that the conditions in lemma 2 hold, $\mathbf{P}$ is invertible and $\gamma<1$.
-Then the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 27 converges to the unique solution of 20.
+Then, the sequence $\{\mathbf{f}^{k}\}$ generated by the iterative scheme 27 converges to the unique solution of 20.
 Proof.
-The proof of theorem 2 can be immediately used here, so we omit the detail. ∎
+The proof of theorem 2 can be directly used here, so we omit the detail. ∎
+Remark 9.
+For the iterative scheme 29, the estimate of section 4.2 presents a linear convergence in the sense of $\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}$.
+Furthermore, by section 4.2, we have that
+$\sigma_{\min}\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\|\mathbf{f}^{k+1}-\mathbf{f}^{*}\|_{2}\leq$ $\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k+1}-\mathbf{f}^{*})\|_{2}$
+$\leq$ $\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\nabla\mathbf{K}(\hat{\mathbf{f}})(\mathbf{f}^{k}-\mathbf{f}^{*})\|_{2}$
+$\leq$ $\sigma_{\max}\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\mathbf{f}^{k}-\mathbf{f}^{*}\|_{2},$
+which implies
+$\|\mathbf{f}^{k+1}-\mathbf{f}^{*}\|_{2}\leq\kappa\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\|\mathbf{f}^{k}-\mathbf{f}^{*}\|_{2}.$ (37)
+Here $\kappa(\cdot)=\sigma_{\max}(\cdot)/\sigma_{\min}(\cdot)$ denotes the condition number.
+Hence, if $\kappa\bigl{(}\nabla\mathbf{K}(\hat{\mathbf{f}})\bigr{)}\bigl{(}\gamma+\zeta(1+\gamma)\bigr{)}\in(0,1)$, the Q-linear convergence of the proposed algorithm is established.
+Here the prefix “Q” stands for “quotient”.
+Remark 10.
+For the iterative scheme 29, a more precise implementation of $\mathbf{P}^{+}$ can generate a smaller residual $\boldsymbol{\delta}^{k}$ in 33.
+Then, there is a smaller $\zeta$ satisfying the condition of 34.
+From 37, it is easy to know that the smaller the $\zeta$, the faster the convergence rate of the algorithm.
 5 Numerical experiments
-We carry out numerical experiments to validate the effectiveness and numerical convergence of the proposed algorithm AFIRE for solving the image reconstruction problem in the geometric-inconsistency DECT in the case of mildly full scan.
-The compared algorithms include some existing methods such as the nonconvex primal-dual algorithm (NCPD) [8], the nonlinear Kaczmarz method using cyclic strategy (NKM) [14], and the iterative FBP reconstruction method (IFBP) [24].
-Moreover, we consider a practical algorithm that uses linear interpolation firstly for the inconsistent data to obtain the data with consistency, and then employ the two-step DDD method to invert the obtained data, where the method for solving the nonlinear systems in the first step is chosen as Newton’s method.
+We carry out numerical experiments to validate the effectiveness and numerical convergence of the proposed algorithm AFIRE for solving the image reconstruction problem in the geometric-inconsistent DECT in the case of mildly full scan.
+The compared algorithms include some existing methods such as the nonconvex primal-dual (NCPD) algorithm [10], the nonlinear Kaczmarz method (NKM) using cyclic strategy [16], and the IFBP reconstruction method [26].
+Moreover, we consider a practical algorithm that firstly processes the geometric-inconsistent data by using linear interpolation to obtain the geometric-consistent data with errors.
+Then the two-step DDD method is used to conduct the reconstruction.
+In the first step, the Newton’s method can be used to estimate the basis sinograms by solving a lot of very small-scale independent nonlinear systems along different X-ray paths.
+In the second step, to reconstruct the basis images, the FBP method is applied to invert the finally obtained basis sinograms.
+To show the intermediate results, the FBP method is also utilized to reconstruct the basis images from the estimated basis sinograms at each Newton iteration in the first step.
 We refer to this algorithm as INTRPL for short.
@@
 The MAC of the basis materials, namely, water and bone, are sourced from [22] and used as the decomposition coefficients $\{\mathbf{b}_{d}\}_{d=1}^{D}$.
+Notably, the use of two basis materials is a standard and physically well-justified practice in clinical DECT.
 Additionally, we employ two full-scan parallel-beam inconsistent geometries for low and high energy spectra, respectively.
@@
 $\Delta_{\mathbf{f}}^{k}:=\frac{\|\mathbf{f}^{k+1}-\mathbf{f}^{k}\|_{2}}{\|\mathbf{f}^{k}\|_{2}},\quad\Delta_{\mathbf{g}}^{k}:=\frac{\|\mathbf{K}(\mathbf{f}^{k+1})-\mathbf{K}(\mathbf{f}^{k})\|_{2}}{\|\mathbf{g}\|_{2}}.$
-In algorithm 3 of AFIRE, as stated in remarks 3 and 4, we take $\hat{\mathbf{f}}=\boldsymbol{0}$, and implement the $\mathbf{P}_{q}^{+}$, $q=1,\ldots,Q$ by FBP method where the 1-bandwidth Ram–Lak filter is used.
+In algorithm 2 of AFIRE, as stated in remarks 6 and 8, we take $\hat{\mathbf{f}}=\boldsymbol{0}$, and implement the $\mathbf{P}_{q}^{+}$, $q=1,\ldots,Q$ by FBP method where the 1-bandwidth Ram–Lak filter is used.
 Moreover, INTRPL and IFBP also use the same FBP method, and NCPD is used to solve a total variation (TV)-regularized nonlinear least square problem with nonnegative constraint.
-For NCPD, the regularization parameter is selected as 00 and $10^{-6}$ for the noiseless and noisy data, respectively.
+For NCPD, the regularization parameter is selected as 0 and $10^{-6}$ for the noiseless and noisy data, respectively.
 Without loss of generality, the initial point for all tested algorithms are set to the zero vector.
 5.2 Simulated noiseless data for the forbild phantom
-In this test, a forbild phantom is composed of the truth water and bone basis images which are both discretized as $128\!\times\!128$ pixels on the square area $[-5,5]\!\times\![-5,5]$ $\text{cm}^{2}$ and take gray values over $[0,1]$, as shown in row 1 of fig.
+In this test, a forbild phantom is composed of the truth water and bone basis images which are both discretized as $\!\times\!128$ 128 pixels on the square area $\!\times\![-5,5]$ [ - 5 , 5 ] $\text{cm}^{2}$ and take gray values over [0,1][0,1][ 0 , 1 ], as shown in row 1 of fig.
 5.
@@
 1 are completely inconsistent.
-Specifically, for the 80-kV spectrum, we utilize 384 parallel projections uniformly sampled on $[-7.05,7.05]$ cm for each of the 384 views, where these views are uniformly distributed over the interval $[0,\pi)$ and can be seen as a mildly full scan.
+Specifically, for the 80-kV spectrum, we utilize 384 parallel projections uniformly sampled on [7.05,7.05][-7.05,7.05][ - 7.05 , 7.05 ] cm for each of the 384 views, where these views are uniformly distributed over the interval $[0,\pi)$ and can be seen as a mildly full scan.
 Under the filtered 140-kV spectrum, the scanning also has the same number of projections and views, but the views are uniformly distributed over the interval $[\pi/768,\pi/768+\pi)$.
@@
 The results also confirm the convergence theory established in theorem 2.
-In contrast, NCPD and NKM are still far from convergence, and INTRPL can only produce low-accuracy result.
+In contrast, NCPD and NKM are still far from convergence, and INTRPL can only produce a low-accuracy result.
 Particularly, it is observed that IFBP does not converge.
+This is because IFBP, as a heuristic approach that designed for geometric-inconsistent case, lacks rigorous theoretical guarantees.
 The metrics of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ over time are also presented in fig.
@@
 5.3 Simulated noisy data for the realistic torso image
-For this test, a real patient torso image is decomposed into the truth water and bone basis images which are both discretized as $256\!\times\!256$ pixels on the square area $[-5,5]\!\times\![-5,5]$ $\text{cm}^{2}$ and take gray values over $[0,1]$, as shown in row 1 of fig.
+For this test, a real patient torso image is decomposed into the truth water and bone basis images which are both discretized as $\!\times\!256$ 256 pixels on the square area $\!\times\![-5,5]$ [ - 5 , 5 ] $\text{cm}^{2}$ and take gray values over [0,1][0,1][ 0 , 1 ], as shown in row 1 of fig.
 9.
@@
 1 are also completely inconsistent.
-More precisely, under the 80-kV spectrum, 768 parallel-beam projections uniformly distributed on $[-7.05,7.05]$ cm are taken for each of 768 views, where these views are evenly sampled on the interval $[0,\pi)$.
+More precisely, under the 80-kV spectrum, 768 parallel-beam projections uniformly distributed on [7.05,7.05][-7.05,7.05][ - 7.05 , 7.05 ] cm are taken for each of 768 views, where these views are evenly sampled on the interval $[0,\pi)$.
 As to the filtered 140-kV spectrum, the same numbers of projections and views are utilized, but the views are uniformly distributed over the interval $[\pi/1536,\pi/1536+\pi)$.
@@
 6 within 100 iterations.
-It is obvious that the metrics of relative errors for the proposed algorithm AFIRE exhibit numerical convergence and relatively high accuracy in less than $15$ iterations, demonstrating its effectiveness and stability for addressing noisy data.
-On the contrary, NCPD and NKM have not converged yet; INTRPL can only produce low-accuracy result; IFBP is divergent after a couple of iterations.
+It is obvious that the metrics of relative errors for the proposed algorithm AFIRE exhibit numerical convergence and relatively high accuracy in less than 151515 iterations, demonstrating its effectiveness and stability for addressing noisy data.
+On the contrary, NCPD and NKM have not converged yet; INTRPL can only produce a low-accuracy result; IFBP is divergent after a couple of iterations.
 […]
@@
 […]
-We also present the curves of metrics $\Delta_{\mathbf{f}}^{k}$ and $\Delta_{\mathbf{g}}^{k}$ in fig. 8, which definitely validate the numerical convergence of the proposed algorithm AFIRE.
+We also present the curves of metrics $\Delta_{\mathbf{f}}^{k}$ and $\Delta_{\mathbf{g}}^{k}$ in fig. 8, which validate the numerical convergence of the proposed algorithm AFIRE.
 Figure 8: Metrics $\Delta_{\mathbf{f}}^{k}$ (left) and $\Delta_{\mathbf{g}}^{k}$ (right) are plotted in semi-log scale as functions of iteration numbers for reconstructing basis images of the real torso image by the proposed algorithm AFIRE from the noisy data.
@@
 Actually, this algorithm can be seen as a framework for solving the image reconstruction problem in MSCT, since there are two main adjustable modules, for instance, including the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$ in algorithm 2 of AFIRE.
-On the other hand, although this algorithm was originally designed to handle the reconstruction problem of geometric inconsistency between low and high X-ray energy spectra, it can be directly used to solve the problem in geometric-consistency situation.
+On the other hand, although this algorithm was originally designed to handle the reconstruction problem of geometric inconsistency between low and high X-ray energy spectra, it can be directly used to solve the problem in geometric-consistent situation.
 Hence, the flexibility and extensibility may enhance the applicability of the proposed algorithm.
@@
 11, 12 and 14.
-All the basis images are discretized as $362\!\times\!362$ pixels on the square area $[-15,15]\!\times\![-15,15]$ $\text{cm}^{2}$.
-In the discussion of the two adjustable modules, under the 80-kV spectrum, we employ 1086 parallel projections evenly sampled on $[-21.15,21.15]$ cm for each of the 900 views, where these views are uniformly distributed over the interval $[0,\pi)$.
+All the basis images are discretized as $\!\times\!362$ 362 pixels on the square area $\!\times\![-15,15]$ [ - 15 , 15 ] $\text{cm}^{2}$.
+In the discussion of the two adjustable modules, under the 80-kV spectrum, we employ 1086 parallel projections evenly sampled on [21.15,21.15][-21.15,21.15][ - 21.15 , 21.15 ] cm for each of the 900 views, where these views are uniformly distributed over the interval $[0,\pi)$.
 Under the filtered 140-kV spectrum, the scanning uses the same numbers of projections and views, but the views are uniformly distributed over the interval $[\pi/1800,\pi/1800+\pi)$.
@@
 Based on our insights in section 3, in the discrete (continuous) case, obviously, for the real constant set $\{C_{d}^{[q]}=0\}_{d,q=1}^{D,Q}$ ($\{C_{d}=0\}_{d=1}^{D}$), there exists $\hat{\mathbf{f}}=\boldsymbol{0}$ ($\hat{\boldsymbol{f}}=\boldsymbol{0}$) fulfilling the condition 23 (the condition 12).
-As we can imagine, those are almost the most simple examples satisfying the required conditions.
+As we can imagine, those are almost the simplest examples that satisfying the required conditions.
 The results in section 5 have demonstrated the effectiveness and numerical convergence of the proposed algorithm when we set each $C_{d}^{[q]}=0$ with $\hat{\mathbf{f}}=\boldsymbol{0}$.
@@
 Additionally, all the $\mathbf{P}_{q}^{+}$, $q=1,\ldots,Q$, are implemented by the FBP with 1-bandwidth Ram–Lak filter.
-As shown in fig. 10, we plot the curves of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in $100$ iterations.
+As shown in fig. 10, we plot the curves of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ in 100100100 iterations.
 […]
@@
 Besides that, the proposed algorithm AFIRE also allows us to adopt different implementations for $\mathbf{P}^{+}$.
-Specifically, we further consider the other two different implementation methods: conjugate gradient method with 20 inner iterations (CG) and limited-memory BFGS method with 60 inner iterations (L-BFGS).
+Specifically, we further consider the other two different implementation methods: conjugate gradient (CG) method with 20 inner iterations and limited-memory BFGS (L-BFGS) method with 60 inner iterations.
 Note that, the numbers of the inner iterations are not fixed, which can be modified on demand.
@@
 From top to bottom: the truths, the results after 100 iterations by the proposed algorithm AFIRE with the implementations of FBP, CG and L-BFGS for the approximated inverse $\mathbf{P}^{+}$.
-Moreover, as indicated in fig. 13, the metric curves of $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ also reveal the similar convergence behavior among those implementations.
 […]
-The implementation that requires inner iterations such as CG and L-BFGS achieves slightly lower error of $\text{RE}_{\mathbf{f}}^{k}$ than the used FBP.
-On the contrary, as evidenced by the comparison of computing time in table 1, FBP is faster than the other two implementations in single (outer) iteration and when the errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ reach $10^{-5}$.
-As proved in theorem 2, more precise implementation of $\mathbf{P}^{+}$ may improve convergence rate of the entire algorithm.
-However, it may increase the computational cost and reduce the efficiency.
+Moreover, as indicated in fig.
+13, the metric curves of $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ reveal the similar convergence behavior among those implementations.
+The implementation that requires inner iterations, such as CG and L-BFGS, achieves a slightly lower $\text{RE}_{\mathbf{f}}^{k}$ than FBP after the same number of outer iterations.
+In contrast, as evidenced by the comparison of running time in table 1, the FBP implementation is faster in a single (outer) iteration and when the errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ reach $10^{-5}$.
+As described in remark 10, a more precise implementation of $\mathbf{P}^{+}$ may improve the convergence rate.
+However, it may increase the computational cost or reduce the efficiency.
 Therefore, there is a trade-off between the reconstructed accuracy and computational efficiency.
@@
   Time (s) FBP CG L-BFGS
-  Single (outer) iteration $0.47$ 0.80 3.22
-$\text{RE}_{\mathbf{f}}^{k}$ reaches $10^{-5}$ $13.99$ 20.70 86.81
-$\text{RE}_{\mathbf{g}}^{k}$ reaches $10^{-5}$ $8.86$ 15.13 61.09
-6.3 Extension to the geometric-consistency situation
-From the derivation in section 3, we know that the proposed algorithm AFIRE can be naturally extended to the geometric-consistency case.
+  Single (outer) iteration 0.470.470.47 0.80 3.22
+$\text{RE}_{\mathbf{f}}^{k}$ reaches $10^{-5}$ 13.9913.9913.99 20.70 86.81
+$\text{RE}_{\mathbf{g}}^{k}$ reaches $10^{-5}$ 8.868.868.86 15.13 61.09
+6.3 Extension to the geometric-consistent situation
+From the derivation in section 3, we know that the proposed algorithm AFIRE can be naturally extended to the geometric-consistent case.
 To illustrate this, what is different from before is that the identical equidistant fan-beam geometry is used for both 80-kV and 140-kV spectra.
@@
 The proposed algorithm is used to reconstruct the basis images from the noiseless data.
-In the geometric-consistency case, the two-step DDD method can be used as a benchmark algorithm for comparison.
-In the first step, to estimate the basis sinograms, we use Newton’s method to solve the associated nonlinear systems in data domain, where the iteration number is 10 and the relative error achieves $10^{-14}$.
-In the second step, to reconstruct the basis images, the FBP with 1-bandwidth Ram–Lak filter is also applied to invert the obtained basis sinograms.
+In the geometric-consistent case, the two-step DDD method can be used as a benchmark algorithm for comparison.
+In the first step, to estimate the basis sinograms in data domain, we use Newton’s method to solve a lot of very small-scale independent nonlinear systems along different X-ray paths.
+When the iteration number is 101010, the relative error has arrived at $10^{-14}$.
+In the second step, to reconstruct the basis images, the FBP with 1-bandwidth Ram–Lak filter is applied to invert the finally obtained basis sinograms.
+To show the intermediate results, the FBP method is also utilized to reconstruct the basis images from the estimated basis sinograms at each Newton iteration in the first step.
 After running algorithm 2 for 100 iterations with setting $\hat{\mathbf{f}}=\boldsymbol{0}$, the obtained basis images and VMIs are shown in fig.
@@
 These comparisons demonstrate that the performance of the proposed algorithm far exceeds the benchmark two-step DDD method.
-Hence, the proposed algorithm cannot only be successfully extended to deal with the geometric-consistency case of MSCT image reconstruction, but also achieve very high accuracy and efficiency.
+Hence, the proposed algorithm cannot only be successfully extended to deal with the geometric-consistent case of MSCT image reconstruction, but also achieve very high accuracy and efficiency.
 Figure 14: From left to right: basis images of water and bone, VMIs at energies 60 keV and 100 keV of a clinical image.
@@
 […]
+6.4 Validation with the basis images generated by randomly sampled Gaussian distribution
+In this work, the AFIRE algorithm is developed for the geometric-inconsistent multispectral CT with a mildly full scan.
+In this case, the number of measurements is comparable to the total size of the employed basis images.
+The sufficient projection data that obtained in this case allows us to treat the basis images as entirely unknown variables, and thus can achieve an efficient reconstruction without priors on the basis images.
+To further validate the statement above, we conduct additional experiments using the basis images generated by randomly sampled Gaussian distribution as truths for the water and bone basis images.
+Specifically, we generate 1000 random array pairs as the basis images using numpy.random with different seeds.
+All the other experimental settings are consistent with those described in section 5.2.
+After running the AFIRE algorithm for 100 iterations, we plot the scattered distributions of the metrics of relative errors $\text{RE}_{\mathbf{f}}^{k}$ and $\text{RE}_{\mathbf{g}}^{k}$ compared to those different random truths in fig.
+16.
+It demonstrates that the AFIRE algorithm can indeed arrive at a high-precision reconstruction even for the randomly generated images.
+This highlights the robustness and effectiveness of the proposed AFIRE algorithm.
+[…]
 7 Conclusion
 In this work, the image reconstruction algorithm was investigated for the nonlinear MSCT when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched.
-We studied the associated CC (DD)-data model, and discovered that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix.
+We studied the associated CC (resp.
+DD)-data model, and discovered that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix.
 Based on the insights, we proposed an accurate and fast image reconstruction algorithm called AFIRE respectively from the continuous, discrete and actual-use perspectives by leveraging the simplified Newton method.
 Under proper conditions, we further proved the convergence of the proposed algorithm AFIRE.
-Furthermore, several numerical experiments were carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in geometric-inconsistency DECT for noiseless and noisy projection data in the case of mildly full scan.
+Furthermore, several numerical experiments were carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in geometric-inconsistent DECT for noiseless and noisy projection data in the case of mildly full scan.
 Through numerical comparisons, it was shown that the proposed algorithm significantly outperforms some existing methods such as NCPD, NKM, IFBP and INTRPL with respect to the accuracy and efficiency.
 We also discussed the flexibility and extensibility of the proposed algorithm.
-It was demonstrated that the proposed algorithm is robust and flexible with regard to the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$.
-Moreover, for the geometric-consistency situation, it was shown that the proposed algorithm has significant advantages over the benchmark two-step DDD method.
+It was demonstrated that the proposed algorithm is robust and flexible with regard to the selection of the special point $\hat{\mathbf{f}}$ and the implementation of the approximated inverse $\mathbf{P}^{+}$, and also can indeed arrive at a high-precision reconstruction even for the randomly generated images.
+Moreover, for the geometric-consistent situation, it was shown that the proposed algorithm has significant advantages over the benchmark two-step DDD method.
 To the best of our knowledge, when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched in MSCT, this is the first literature that uses the characteristics of the derivative operator (gradient) of the associated nonlinear mapping to construct the algorithm for accurate and fast image reconstruction.
@@
 Proof.
-For any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$, $K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})$ can be rewritten as the function $H_{q}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ which is given by
+For any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q}$, $K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})$ can be rewritten as the function $H_{q}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ which is given by
 $\boldsymbol{z}$ $\longmapsto\ln\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)\mathrm{d}E,$
-where $\boldsymbol{z}=[z_{1},\ldots,z_{D}]^{\boldsymbol{\mathsf{T}}}$ and $z_{d}=\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})$. It is easy to compute that
-$\frac{\partial H_{q}(\boldsymbol{z})}{\partial z_{d}}=-\int_{0}^{E_{\max}}\frac{s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)}{\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)\mathrm{d}E}b_{d}(E)\mathrm{d}E.$ (33)
-By 33 and the nonnegative bounded property of functions $\{b_{d}\}_{d=1}^{D}$, we have
+where $\boldsymbol{z}=[z_{1},\ldots,z_{D}]^{\boldsymbol{\mathsf{T}}}$ and $z_{d}=\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})$. It is easy to compute that
+$\frac{\partial H_{q}(\boldsymbol{z})}{\partial z_{d}}=-\int_{0}^{E_{\max}}\frac{s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)}{\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)z_{d}\right)\mathrm{d}E}b_{d}(E)~\mathrm{d}E.$ (38)
+By 38 and the nonnegative bounded property of functions $\{b_{d}\}_{d=1}^{D}$, we have
 $\|\nabla H_{q}(\boldsymbol{z})\|_{1}=\sum_{d=1}^{D}\left|\frac{\partial H_{q}(\boldsymbol{z})}{\partial z_{d}}\right|\leq\sum_{d=1}^{D}\sup_{E}b_{d}(E).$ (39)
-Notice that $K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})=0$ for any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$, and using 34, we obtain
-$\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\left|H_{q}(\boldsymbol{z})-H_{q}(0)\right|\\ \leq\left(\sup_{\boldsymbol{z}}\|\nabla H_{q}(\boldsymbol{z})\|_{1}\right)\|\boldsymbol{z}\|_{1}\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|.$ (35)
-Consequently, for any $\boldsymbol{f}\in\mathcal{F}^{D}$, by 35, we have
-$\|K^{[q]}(\boldsymbol{f})\|_{\mathcal{G}}$ $=\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
-$\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\Bigl{(}\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|\Bigr{)}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
-$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\sum_{d=1}^{D}\left|\mathcal{P}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
-$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\|\mathcal{P}f_{d}\|_{\mathcal{G}},$
+Notice that $K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})=0$ for any $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q}$, and using 39, we obtain
+$\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-K^{[q]}(\boldsymbol{0})(\boldsymbol{\omega},\boldsymbol{x})\big{|}=\left|H_{q}(\boldsymbol{z})-H_{q}(0)\right|\\ \leq\left(\sup_{\boldsymbol{z}}\|\nabla H_{q}(\boldsymbol{z})\|_{1}\right)\|\boldsymbol{z}\|_{1}\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\left|\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|.$ (40)
+Consequently, for any $\boldsymbol{f}\in\mathbfcal{F}^{D}$, by 40, we have
+$\|K^{[q]}(\boldsymbol{f})\|_{\mathcal{G}_{q}}$ $=\left(\int_{\mathbb{S}^{n-1}_{q}}\int_{\Omega_{q}}\big{|}K^{[q]}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})\big{|}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
+$\leq\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}_{q}}\int_{\Omega_{q}}\Bigl{(}\sum_{d=1}^{D}\left|\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|\Bigr{)}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
+$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\left(\int_{\mathbb{S}^{n-1}_{q}}\int_{\Omega_{q}}\sum_{d=1}^{D}\left|\mathcal{P}_{q}f_{d}(\boldsymbol{\omega},\boldsymbol{x})\right|^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}\right)^{1/2}$
+$\leq\sqrt{D}\left(\sum_{d=1}^{D}\sup_{E}b_{d}(E)\right)\sum_{d=1}^{D}\|\mathcal{P}_{q}f_{d}\|_{\mathcal{G}_{q}},$
 where $\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}$ denotes the Euclidean measure on the hyperplane $\boldsymbol{\omega}^{\perp}$.
-Since $\mathcal{P}f_{d}\in\mathcal{G}$ for each $d$, we obtain $K^{[q]}(\boldsymbol{f})\in\mathcal{G}$.
-For any $\boldsymbol{h}=[h_{1},\ldots,h_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathcal{F}^{D}$, by the chain rule, we compute
-$\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}h_{d}\right)\mathrm{d}E,$ (36)
+Since $\mathcal{P}_{q}f_{d}\in\mathcal{G}_{q}$ for each $d$, we obtain $K^{[q]}(\boldsymbol{f})\in\mathcal{G}_{q}$.
+For any $\boldsymbol{h}=[h_{1},\ldots,h_{D}]\in\mathbfcal{F}^{D}$, by the chain rule, we compute
+$\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}h_{d}\right)\mathrm{d}E,$ (41)
 where $\Phi^{[q]}(\boldsymbol{f};E)$ is given as in 11 such that
-$\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathrm{d}E=1\quad\text{for}~{}q=1,\ldots,Q.$
-Then, from 36, we have the following estimate
-$\big{\|}\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}\big{\|}_{\mathcal{G}}$ $=\big{\|}\sum_{d=1}^{D}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)b_{d}(E)\mathcal{P}h_{d}\mathrm{d}E\big{\|}_{\mathcal{G}}$ (37)
-$\leq\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\big{\|}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathcal{P}h_{d}\mathrm{d}E\big{\|}_{\mathcal{G}}$
-$=\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\left\|\mathcal{P}h_{d}\right\|_{\mathcal{G}}$
-$\leq\left(\sup_{d,E}b_{d}(E)\right)C\sum_{d=1}^{D}\|h_{d}\|_{\mathcal{F}}.$
-Note that the last inequality follows readily from the result in [30, theorem 2.10], and the $C$ is a positive constant depending only on the $\alpha=-1/2$ and the domain $\Omega$.
-Hence, using 36 and 37, we conclude that $K^{[q]}$ is Gâteaux differentiable, and its Gâteaux derivative is computed as 10.
-Furthermore, from the continuity of $\mathcal{P}$, it is easy to obtain that the mapping $\boldsymbol{f}\mapsto\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$ is continuous.
-By [12, proposition 3.2.15], $K^{[q]}$ is Fréchet differentiable and its Fréchet derivative is equal to $\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$.
+$\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)~\mathrm{d}E=1\quad\text{for}~q=1,\ldots,Q.$
+Then, from 41, we have the following estimate
+$\big{\|}\partial_{\varepsilon}K^{[q]}(\boldsymbol{f}+\varepsilon\boldsymbol{h})\mid_{\varepsilon=0}\big{\|}_{\mathcal{G}_{q}}$ $=\big{\|}\sum_{d=1}^{D}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)b_{d}(E)\mathcal{P}_{q}h_{d}~\mathrm{d}E\big{\|}_{\mathcal{G}_{q}}$
+$\leq\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\big{\|}\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\mathcal{P}_{q}h_{d}~\mathrm{d}E\big{\|}_{\mathcal{G}_{q}}$
+$=\left(\sup_{d,E}b_{d}(E)\right)\sum_{d=1}^{D}\left\|\mathcal{P}_{q}h_{d}\right\|_{\mathcal{G}_{q}}$
+$\leq\left(\sup_{d,E}b_{d}(E)\right)C\sum_{d=1}^{D}\|h_{d}\|_{\mathcal{F}}.$ (42)
+Note that the last inequality follows readily from the result in [32, theorem 2.10], and the $C$ is a positive constant depending only on $\alpha=-1/2$ and the domain $\Omega$.
+Hence, using 41 and section A.1, we conclude that $K^{[q]}$ is Gâteaux differentiable, and its Gâteaux derivative is computed as 10.
+Furthermore, from the continuity of $\mathcal{P}_{q}$, it is easy to obtain that the mapping $\boldsymbol{f}\mapsto\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$ is continuous.
+By [14, proposition 3.2.15], $K^{[q]}$ is Fréchet differentiable and its Fréchet derivative is equal to $\mathcal{D}K^{[q]}(\boldsymbol{f},\cdot)$.
 By the definition of the norm of Cartesian space, we can further prove that $\boldsymbol{K}$ is also Fréchet differentiable.
@@
 where
-$\psi^{[q]}_{d}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})=\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})b_{d}(E)\mathrm{d}E.$
+$\psi^{[q]}_{d}(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})=\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})b_{d}(E)~\mathrm{d}E.$
 Since $\Phi^{[q]}(\boldsymbol{f};\,\cdot\,)$ as given in 11 is a nonnegative normalized function with respect to $E$, together with the given conditions, we have
-$\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\|_{F}^{2}$ $=\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\big{(}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})-\Phi^{[q]}(\hat{\boldsymbol{f}};E)\big{)}b_{d}(E)\mathrm{d}E\right)^{2}$
-$\leq\eta^{2}\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E\right)^{2}=\eta^{2}\|\boldsymbol{\varphi}\|_{F}^{2}.$
-Subsequently, for any $\boldsymbol{z}\in\mathbb{R}^{D}$, by Cauchy–Schwarz inequality and the property of matrix norm, we have
+$\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-$ $\boldsymbol{\varphi}\|_{F}^{2}$
+$=\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\big{(}\Phi^{[q]}(\boldsymbol{f};E)(\boldsymbol{\omega},\boldsymbol{x})-\Phi^{[q]}(\hat{\boldsymbol{f}};E)\big{)}b_{d}(E)~\mathrm{d}E\right)^{2}$
+$\leq\eta^{2}\sum_{q=1}^{Q}\sum_{d=1}^{D}\left(\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)~\mathrm{d}E\right)^{2}=\eta^{2}\|\boldsymbol{\varphi}\|_{F}^{2}.$
+Subsequently, for any $\boldsymbol{z}\in\mathbb{R}^{D}$, by Cauchy–Schwarz inequality and the property of the matrix norm, we have
 $\|(\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi})\boldsymbol{z}\|_{2}$ $\leq\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\|_{2}\|\boldsymbol{z}\|_{2}\leq\|\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\|_{F}\|\boldsymbol{z}\|_{2}$
 $\leq\eta\|\boldsymbol{\varphi}\|_{F}\|\boldsymbol{z}\|_{2}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\boldsymbol{\varphi}\boldsymbol{z}\|_{2}.$
-Hence, for any $\boldsymbol{f},\boldsymbol{h}\in\mathcal{F}^{D}$, based on the formula of Fréchet derivative of $\boldsymbol{K}$ and the above estimation, we obtain
-$\big{\|}\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\big{\|}_{\mathcal{G}^{Q}}^{2}$ $=\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\left\|\left(\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\right)\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})\right\|_{2}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}$
+Hence, for any $\boldsymbol{f},\boldsymbol{h}\in\mathbfcal{F}^{D}$, based on the formula of Fréchet derivative of $\boldsymbol{K}$ and the above estimation, we obtain
+$\big{\|}\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},$ $\boldsymbol{h})\big{\|}_{\mathbfcal{G}^{Q}}^{2}$
+$=\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}\left\|\left(\Psi(\boldsymbol{f})(\boldsymbol{\omega},\boldsymbol{x})-\boldsymbol{\varphi}\right)\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})\right\|_{2}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}$
 $\leq\int_{\mathbb{S}^{n-1}}\int_{\boldsymbol{\omega}^{\perp}}(\eta\kappa_{F}(\boldsymbol{\varphi}))^{2}\|\boldsymbol{\varphi}\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})\|_{2}^{2}\mathrm{d}\boldsymbol{\mu}_{\boldsymbol{x}}\mathrm{d}\boldsymbol{\omega}$
-$=(\eta\kappa_{F}(\boldsymbol{\varphi}))^{2}\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}^{2},$
-where $\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})=[\mathcal{P}h_{1}(\boldsymbol{\omega},\boldsymbol{x}),\ldots,\mathcal{P}h_{D}(\boldsymbol{\omega},\boldsymbol{x})]^{\boldsymbol{\mathsf{T}}}\in\mathbb{R}^{D}$.
+$=(\eta\kappa_{F}(\boldsymbol{\varphi}))^{2}\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathbfcal{G}^{Q}}^{2},$
+where $\boldsymbol{\mathcal{P}}\boldsymbol{h}(\boldsymbol{\omega},\boldsymbol{x})=[\mathcal{P}_{1}h_{1}(\boldsymbol{\omega},\boldsymbol{x}),\ldots,\mathcal{P}_{D}h_{D}(\boldsymbol{\omega},\boldsymbol{x})]^{\boldsymbol{\mathsf{T}}}\in\mathbb{R}^{D}$.
 The desired result is obtained.